### v2

# Diversity or Precision? A Deep Dive into Next Token Prediction

Abstract
Recent advancements have shown that reinforcement learning (RL) can substantially improve the reasoning abilities of large language models (LLMs). The effectiveness of such RL training, however, depends critically on the exploration space defined by the pre-trained model’s token-output distribution. In this paper, we revisit the standard cross-entropy loss, interpreting it as a specific instance of policy gradient optimization applied within a single-step episode. To systematically study how the pre-trained distribution shapes the exploration potential for subsequent RL, we propose a generalized pre-training objective that adapts on-policy RL principles to supervised learning. By framing next-token prediction as a stochastic decision process, we introduce a reward-shaping strategy that explicitly balances diversity and precision. Our method employs a positive reward scaling factor to control probability concentration on ground-truth tokens and a rank-aware mechanism that treats high-ranking and low-ranking negative tokens asymmetrically. This allows us to reshape the pre-trained token-output distribution and investigate how to provide a more favorable exploration space for RL, ultimately enhancing end-to-end reasoning performance. Contrary to the intuition that higher distribution entropy facilitates effective exploration, we find that imposing a precision-oriented prior yields a superior exploration space for RL.
footnotetext: Project Lead.
1 Introduction
Recent advancements have demonstrated that reinforcement learning (RL) (Bai et al., 2022; Guo et al., 2025) can significantly enhance the reasoning capabilities of large language models (LLMs) (Google DeepMind, 2025; Guo et al., 2025; Anthropic, 2025; Kimi et al., 2025). By utilizing verifiable rewards, such as passing unit tests or deriving correct mathematical solutions, LLMs evolve from merely mimicking human data to actively searching for optimal reasoning paths (Guo et al., 2025). On-policy training paradigms have proven effective in unlocking the potential of pre-trained LLMs, prompting researchers to investigate how token output distributions influence RL. Recent studies (Wang et al., 2025; Zhu et al., 2025b; Cui et al., 2025; Gandhi et al., 2025) indicate that uncertainty in chain-of-thought reasoning is concentrated within a small subset of high-entropy forking tokens that govern pivotal decisions, while the majority of tokens exhibit low entropy. This observation underscores the critical impact of the pre-trained model’s output distribution on subsequent RL outcomes.
Concurrently, researchers have explored next-token and next-segment reasoning objectives to derive self-supervised signals from massive unlabeled pre-training corpora (Zelikman et al., 2024; Dong et al., 2025; Li et al., 2025; Xing et al., 2025). Applying RL to the pre-training corpus suggests a theoretical bridge connecting pre-training and RL. Specifically, next-token prediction can be reformulated as a reasoning task optimized via RL algorithms, where the model receives verifiable rewards for accurately predicting the subsequent token according to a given context. Notably, if the intermediate reasoning process is omitted, resulting in the direct generation of the answer, this procedure becomes analogous to standard pre-training. From the perspective of policy optimization, next-token prediction serves a foundational role by defining the initial policy distribution for subsequent RL. This distribution establishes the model’s behavioral trajectory and implicitly constrains its exploration space, thereby determining which reasoning paths the model prioritizes during RL.
Motivated by this connection, we revisit the cross-entropy loss for next token prediction. Although traditionally viewed as a supervised metric, cross-entropy can be interpreted as a specific instance of policy gradient optimization within a single-step episode (Wu et al., 2025; Ming et al., 2025). This interpretation suggests that next-token prediction inherently permits an on-policy perspective, even though standard teacher forcing utilizes off-policy samples drawn directly from the training corpus distribution. From an entropy perspective, cross-entropy implicitly assigns maximal reward to the single ground-truth token while uniformly suppressing all negative tokens. Building on this insight, we aim to establish a unified pre-training objective that subsumes cross-entropy as a special case, enabling a systematic study of how reward configurations during pre-training influence subsequent RL dynamics.
In this paper, we propose a generalized objective that integrates on-policy training principles into supervised learning. By formulating next-token prediction as a stochastic decision process, we expose the intrinsic reward mechanism of cross-entropy and introduce a reward-shaping strategy. This approach explicitly regulates the trade-off between diversity and precision during pre-training, rather than deferring this balance to subsequent RL stages. Specifically, we introduce a positive reward scaling factor to control the concentration of probability mass on ground-truth tokens, and we differentiate between high-ranking and low-ranking negative tokens to modulate suppression asymmetrically. This strategy allows us to reshape the token output distribution and systematically analyze the relationship between pre-training objectives and RL exploration. Contrary to the conventional intuition that higher distribution entropy facilitates effective exploration, our findings reveal that imposing a precision-oriented prior yields a superior exploration space for RL, ultimately enhancing end-to-end reasoning performance.
Our main contributions are summarized as follows:
  • We propose a generalized pre-training objective for next-token prediction that incorporates a reward-shaping strategy, utilizing a positive reward scaling factor and rank-aware negative suppression.
  • We investigate how reshaping the token output distribution during pre-training modulates the exploration space for subsequent RL, thereby impacting end-to-end reasoning performance.
  • We demonstrate that a precision-oriented pre-training prior provides a more effective initialization for RL than high-entropy distributions, leading to improved reasoning capabilities.
2 Method
2.1 Next Token Prediction
Autoregressive LLMs are typically trained using a next-token prediction objective. This process can be formulated as a sequential decision-making problem where the LLM functions as a stochastic policy $\pi_{\theta}$.
Let $X=\{x_{1},x_{2},\cdots,x_{n}\}$ denote a sequence of $n$ tokens. At step $t$, the state $s_{t}$ is defined by the prefix $X_{<t}=\{x_{1},x_{2},\cdots,x_{t-1}\}$. The action $a_{t}$ corresponds to the next token, sampled from the vocabulary $V$ according to the policy $\pi_{\theta}(\cdot\mid s_{t})$. The training objective optimizes the parameters $\theta$ to maximize the expected cumulative reward:
$J(\theta)=\mathbbm{E}_{\tau\sim\pi_{\theta}}\Big[\sum_{t=1}^{n}r(s_{t},a_{t})\Big],$ (1)
where $\tau=(s_{1},a_{1},s_{2},a_{2},\cdots)$ represents a trajectory sampled from $\pi_{\theta}$, and $r(s_{t},a_{t})$ is the scalar reward received for taking action $a_{t}$ in state $s_{t}$. The policy gradient can be derived as:
$\nabla_{\theta}J(\theta)=\mathbbm{E}_{\tau\sim\pi_{\theta}}\Big[\sum_{t=1}^{n}R(\tau)\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})\Big],$ (2)
where $R(\tau)=\sum_{t^{\prime}=1}^{n}r(s_{t^{\prime}},a_{t^{\prime}})$. To reduce variance without introducing bias, the total return $R(\tau)$ is typically replaced by the return-to-go $G_{t}=\sum_{t^{\prime}=t}^{n}r(s_{t^{\prime}},a_{t^{\prime}})$, often incorporating a baseline $b(s_{t})$ for variance reduction:
$\nabla_{\theta}J(\theta)=\mathbbm{E}_{\tau\sim\pi_{\theta}}\Big[\sum_{t=1}^{n}(G_{t}-b(s_{t}))\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})\Big].$ (3)
Building upon Equation 3, we treat the generation of a single token as a complete episode (Ming et al., 2025). The objective for a fixed state $s_{t}$ simplifies to:
$J_{t}(\theta\mid s_{t})=\mathbbm{E}_{a_{t}\sim\pi_{\theta}(\cdot\mid s_{t})}[r(s_{t},a_{t})],$ (4)
yielding the gradient:
$\nabla_{\theta}J_{t}(\theta\mid s_{t})=\mathbbm{E}_{a_{t}\sim\pi_{\theta}(\cdot\mid s_{t})}\big[r(s_{t},a_{t})\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})\big].$ (5)
Crucially, for Equation 5 to remain consistent with the cumulative reward structure of Equation 3, the reward $r(s_{t},a_{t})$ must depend solely on the immediate state-action pair.
2.2 Revisiting Cross-Entropy
LLM pre-training is generally cast as a supervised learning process designed to maximize the log-likelihood of the ground-truth token $x_{t}$ given the context $s_{t}=X_{<t}$:
$J_{\text{CE}}(\theta)=\log\pi_{\theta}(x_{t}\mid s_{t}).$ (6)
The gradient of this objective explicitly maximizes the probability of the ground-truth token:
$\nabla_{\theta}J_{\text{CE}}(\theta)=\nabla_{\theta}\log\pi_{\theta}(x_{t}\mid s_{t}).$ (7)
We can express this gradient as an expectation over the full policy distribution $\pi_{\theta}(\cdot\mid s_{t})$, encompassing both positive ($a_{t}=x_{t}$) and negative ($a_{t}\neq x_{t}$) tokens. By invoking the log-derivative identity $\nabla_{\theta}\log\pi_{\theta}(x)=\frac{\nabla_{\theta}\pi_{\theta}(x)}{\pi_{\theta}(x)}$ and introducing the indicator function $\mathbbm{1}(a_{t}=x_{t})$, we expand the gradient into a summation over the vocabulary $V$:
$\nabla_{\theta}J_{\text{CE}}(\theta)$ $=\frac{1}{\pi_{\theta}(x_{t}\mid s_{t})}\nabla_{\theta}\pi_{\theta}(x_{t}\mid s_{t})$
$=\frac{1}{\pi_{\theta}(x_{t}\mid s_{t})}\sum_{a_{t}\in V}\mathbbm{1}(a_{t}=x_{t})\nabla_{\theta}\pi_{\theta}(a_{t}\mid s_{t}).$ (8)
We recover the probability density using the substitution $\nabla_{\theta}\pi_{\theta}(a_{t}\mid s_{t})=\pi_{\theta}(a_{t}\mid s_{t})\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})$, and then form an expectation:
$\nabla_{\theta}J_{\text{CE}}(\theta)$ $=\sum_{a_{t}\in V}\pi_{\theta}(a_{t}\mid s_{t})\left[\frac{\mathbbm{1}(a_{t}=x_{t})}{\pi_{\theta}(a_{t}\mid s_{t})}\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})\right]$
$=\mathbbm{E}_{a_{t}\sim\pi_{\theta}(\cdot\mid s_{t})}\left[r_{\text{CE}}(s_{t},a_{t})\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})\right].$ (9)
In supervised training, the ground-truth token $x_{t}$ is deterministically defined by the dataset. Consequently, the indicator $\mathbbm{1}(a_{t}=x_{t})$ evaluates the action $a_{t}$ against a static property of $s_{t}$, ensuring that the derived intrinsic reward depends exclusively on information available at step $t$. Comparing Equation 9 with Equation 5 reveals the intrinsic reward function of cross-entropy:
$r_{\text{CE}}(s_{t},a_{t})=\text{sg}(\frac{\mathbbm{1}(a_{t}=x_{t})}{\pi_{\theta}(a_{t}\mid s_{t})}),$ (10)
where $\text{sg}(\cdot)$ denotes the stop-gradient operator. Equation 10 demonstrates that when the sampled action matches the ground truth ($a_{t}=x_{t}$), the reward is scaled by the inverse probability $\frac{1}{\pi_{\theta}(x_{t}\mid s_{t})}$. On the contrary, for all negative tokens, the intrinsic reward is exactly 0. Unlike RL scenarios where negative actions are often explicitly penalized, cross-entropy achieves suppression of negative tokens implicitly through the Softmax normalization constraint $\sum\limits_{a_{t}\in V}\pi_{\theta}(a_{t}\mid s_{t})=1$. By increasing the probability of the positive tokens via positive rewards, the probabilities of competing tokens are forced to decrease.
2.3 Diversity or Precision
As derived in Equation 10, the intrinsic reward of the cross-entropy objective implicitly balances diversity and precision. To explicitly regulate the trade-off between these two objectives, we propose a generalized reward function designed to independently control the influence of positive and negative tokens.
First, we introduce a modulating factor to scale the reward associated with the ground-truth token. Let $a_{t}$ denote the generated token and $x_{t}$ the ground truth, we define the modified positive reward as:
$\bar{r}_{\text{pos}}(s_{t},a_{t})=\text{sg}((\frac{1}{\pi_{\theta}(a_{t}\mid s_{t})})^{(1-\pi_{\theta}(a_{t}\mid s_{t}))^{\beta}}),$ (11)
where $(1-\pi_{\theta}(a_{t}\mid s_{t}))^{\beta}$ serves as a positive reward scaling factor. Equation 11 facilitates the control of global entropy. Specifically, when $\beta<0$, the reward is amplified relative to the baseline ($\beta=0$). This produces large gradient updates that aggressively concentrate probability mass onto the ground truth, collapsing the distribution and minimizing global entropy. Conversely, $\beta>0$ attenuates the reward signal. In this regime, the model is less penalized for assigning a lower probability to the ground truth, allowing the policy to maintain a flatter distribution with higher entropy.
Second, while standard cross-entropy assigns zero reward to all negative tokens, we propose shaping the negative distribution to control local entropy. Let $\mathcal{K}_{t}=\text{TopK}(\pi_{\theta}(\cdot\mid s_{t}),k)$ denote the set of the top-$k$ predicted tokens, we define the negative reward as:
$\bar{r}_{\text{neg}}(s_{t},a_{t})=$ $\tilde{\lambda}\cdot\mathbbm{1}(a_{t}\in\mathcal{K}_{t}\land a_{t}\neq x_{t})+\hat{\lambda}\cdot\mathbbm{1}(a_{t}\notin\mathcal{K}_{t}\land a_{t}\neq x_{t}).$ (12)
As shown in Equation 12, we assign a reward $\tilde{\lambda}$ to high-ranking negative tokens to prevent the model from becoming overly confident in the ground truth alone, thereby reserving probability mass for plausible alternatives. Meanwhile, to suppress low-probability tail tokens, we apply a reward $\hat{\lambda}$ to tokens falling outside $\mathcal{K}_{t}$, forcing the distribution to concentrate on the head.
Finally, the generalized reward function for the single-step objective is defined as:
$\bar{r}(s_{t},a_{t})$ $=\bar{r}_{\text{pos}}(s_{t},a_{t})\cdot\mathbbm{1}(a_{t}=x_{t})+\bar{r}_{\text{neg}}(s_{t},a_{t})\cdot\mathbbm{1}(a_{t}\neq x_{t}).$ (13)
Notably, the setting $\beta=0,\tilde{\lambda}=0,\hat{\lambda}=0$ recovers standard cross-entropy.
3 Experiments
3.1 Training Settings
The training pipeline proceeds in three stages: pre-training, mid-training, and RLVR. Adhering to the Qwen3 (Yang et al., 2025), we develop LLMs using both dense and MoE architectures. Specifically, we develop a series of LLMs, which include 1B and 4B dense models, as well as 5B-A0.3B, 10B-A0.5B and 20B-A1B MoE models. Moreover, we conduct the complete training pipeline on the 4B, 10B-A0.5B and 20B-A1B models, while the 1B and 5B-A0.3B models undergo the pre-training stage only. More training details are provided in Appendix A and Appendix B.
Training Data. For pre-training, we curate a corpus of 500B tokens primarily focused on general knowledge. This is followed by a mid-training stage comprising 100B tokens, which incorporates approximately 5% synthetic data and significantly increases the proportion of reasoning-oriented content. Crucially, we deliberately exclude the synthetic long-reasoning data from all training stages to accurately observe the activation trends of the model’s long-CoT reasoning capabilities. The RL stage prioritizes mathematical reasoning tasks, as the emergence of long-reasoning capabilities is typically associated with these domains.
Hyperparameters. Hyperparameters are maintained across the pre-training and mid-training stages. Our goal is to investigate how different reward shaping strategies influence end-to-end performance. Consequently, we perform specific reward configurations for positive tokens ($\beta=-0.25$ and $\beta=0.5$) and negative tokens ($\hat{\lambda}=-0.1,\tilde{\lambda}=0,k=100$ and $\hat{\lambda}=0,\tilde{\lambda}=0.1,k=100$). Employing these distinct hyperparameter configurations allows us to isolate the specific effects of positive and negative reward signals.
3.2 Evaluation Settings
Evaluation of Base Models. Our comprehensive evaluation of base models assesses five core capabilities: general knowledge, logic reasoning, commonsense reasoning, mathematics, and coding. The evaluation is conducted using 19 distinct benchmarks:
  • General Knowledge: MMLU (Hendrycks et al., 2020)(4-shot, CoT), MMLU-Pro (Wang et al., 2024)(5-shot, CoT), TriviaQA (Joshi et al., 2017)(5-shot), and NaturalQuestions (Kwiatkowski et al., 2019)(5-shot).
  • Commonsense Reasoning: Hellaswag (Zellers et al., 2019)(0-shot), SIQA (Sap et al., 2019)(0-shot), PIQA (Bisk et al., 2020)(0-shot), WinoGrande (Sakaguchi et al., 2021)(0-shot), OpenBookQA (Mihaylov et al., 2018)(5-shot), and CommonsenseQA (Talmor et al., 2018)(5-shot)
  • Logic Reasoning: ARC-Easy (Clark et al., 2018)(0-shot), ARC-Challenge (Clark et al., 2018)(0-shot), and BBH (Suzgun et al., 2022)(3-shot, CoT)
  • Mathematics: GSM8K (Cobbe et al., 2021)(4-shot, CoT), MATH-500 (Lightman et al., 2023)(4-shot, CoT), Minerva (Lewkowycz et al., 2022)(4-shot, CoT), and OlympiadBench (He et al., 2024)(0-shot).
  • Coding: HumanEval+ (Liu et al., 2023)(0-shot) and MBPP+ (Liu et al., 2023)(3-shot).
Specifically, general knowledge and commonsense reasoning evaluate the model’s knowledge-base capabilities, whereas logical reasoning, mathematics, and coding probe its reasoning-base capabilities. Moreover, we employ the $\text{Pass@}k$ metric to evaluate the model’s upper-bound capability for tasks requiring mathematical reasoning and code generation. Pass@k measures the probability that at least one correct solution is present within $k$ independent attempts. We utilize the unbiased estimator of $\text{Pass@}k$ (Chen, 2021), which is defined as:
$\text{Pass@}k=1-\frac{\binom{m-c}{k}}{\binom{m}{k}}$, (14)
where $m$ represents the total number of sampled responses generated per prompt, and $c$ denotes the count of correct responses among those $m$ samples. We sample $m=128$ responses with temperature 0.7 and top-p 0.95 and report Pass@64 metric. Notably, we configure the maximum output length to 4K for pre-trained models and 16K for mid-trained models.
Evaluation of RL Models. For RL models evaluation, we employ various mathematics benchmarks, including AMC23 (MAA, b), AIME (MAA, a), MATH-500 (Lightman et al., 2023), Minerva (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). We sample 128 responses per problem and report Avg@128, Cons@128, and Pass@64 metrics. Specifically, $\text{Avg@}128$ represents the average accuracy across all 128 samples, while $\text{Cons@}128$ refers to the majority voting accuracy. Similarly, we configure the maximum output length to 16K for RL models.
Figure 1: Changes of PPL and entropy during pre-training across 1B and 4B dense models, developed based on different configurations.
Figure 2: Changes of PPL and entropy during pre-training across 5B-A0.3B, 10B-A0.5B and 20B-A1B MoE models, developed based on different configurations.
3.3 Pre-Training
Our analysis of the proposed generalized training objective reveals that it effectively regulates the trade-off between diversity and precision by strategically varying reward configurations. As illustrated in Figure 1 and Figure 2, perplexity (PPL) consistently converges to comparable low values across both dense (1B, 4B) and MoE (5B-A0.3B, 10B-A0.5B, 20B-A1B) architectures. This demonstrates that, within a specific range, modifying the reward function modulates training dynamics without compromising final predictive accuracy. The parameter $\beta$ serves as a potent global entropy regulator. Specifically, setting $\beta<0$ significantly reduces entropy, resulting in a more peaked and confident token distribution by amplifying rewards for ground turth tokens. Conversely, $\beta>0$ maintains higher entropy and a flatter distribution, thereby promoting diversity in the generated output. Meanwhile, the parameters $\hat{\lambda}$ and $\tilde{\lambda}$ facilitate local entropy fine-tuning. These parameters shape the token distribution by either rewarding ($\hat{\lambda}=0,\tilde{\lambda}=0.1,k=100$) or penalizing ($\hat{\lambda}=-0.1,\tilde{\lambda}=0,k=100$) negative tokens, enabling granular control over the training process.
Furthermore, we analyze the evolution of model performance during pre-training to investigate the dynamics and specific impact of the proposed reward function. As depicted in Figure 3, larger models consistently achieve substantially higher final performance than smaller models after processing an equivalent number of training tokens. This confirms that explicitly regulating the diversity-precision trade-off is an orthogonal mechanism that does not interfere with the fundamental scaling properties of language models. Crucially, configurations that prioritize lowering global entropy ($\beta<0$) or maintaining high local entropy ($\hat{\lambda}=-0.1,\tilde{\lambda}=0,k=100$) demonstrate superior performance and scaling behavior. Although these settings may not yield optimal initial performance in smaller models, they exhibit enhanced growth potential as model size increases. This suggests that with greater model capacity, strategies that promote precision, either globally via generously rewarding positive tokens or locally by aggressively penalizing tail negative tokens, lead to better performance growth compared to the baseline.
Figure 3: Changes of performance during pre-training across models with various model parameters, developed based on dense and MoE architectures under different configurations.
Figure 4: Changes of performance during mid-training across 4B dense, 10B-A0.5B MoE and 20B-A1B MoE models, developed based on different configurations.
3.4 Mid-Training
Subsequently, we evaluate the evolution of model performance during the mid-training stage, spanning from 0B to 100B tokens. As depicted in Figure 4, the choice of $\beta$ significantly influences training dynamics. We observe a consistent trend where a negative value, specifically $\beta=-0.25$, yields the best results. This configuration consistently outperforms the baseline ($\beta=0$) across both dense and MoE models in knowledge and reasoning tasks. Conversely, a positive setting ($\beta=0.50$) does not demonstrate consistent superior performance comparing to the baseline. Similar to the observations with $\beta$, a slight negative adjustment appears beneficial. The configuration $\hat{\lambda}=-0.1,\tilde{\lambda}=0,k=100$ generally matches or slightly surpasses the performance of the standard CE baseline. However, when shifting to $\hat{\lambda}=0.1,\tilde{\lambda}=0,k=100$, performance exhibited uncertainty in knowledge-intensive scenarios. In reasoning tasks, the performance remained comparable to the standard CE baseline.
Figure 5: Changes of performance during RL training across various actor models, developed based on a 4B dense architecture under different configurations.
Figure 6: Changes of performance during RL training across various actor models, developed based on a 10B-A0.5B MoE architecture under different configurations.
Figure 7: Changes of performance during RL training across various actor models, developed based on a 20B-A1B MoE architecture under different configurations.
3.5 Reinforcement Learning
Finally, we investigate the performance dynamics during the RL training stage across various actor models, as illustrated in Figures 5 to 7. Pre-trained models derived from different reward configurations exhibit distinct output distributions, leading to significant variations in subsequent RL and end-to-end reasoning performance. Regarding the global entropy regulator $\beta$, we observe a consistent and robust trend across both the 4B dense and 10B-A0.5B MoE models. Specifically, the global low entropy setting ($\beta=-0.25$) yields superior performance trajectories. This configuration consistently outperforms the global high entropy setting across all evaluated metrics, including Avg@128, Cons@128, and Pass@64. Furthermore, the configuration $\hat{\lambda}=-0.1,\tilde{\lambda}=0,k=100$ demonstrates a significant advantage, consistently achieving the highest performance and notably surpassing the baseline on the 10B-A0.5B MoE model. For the 4B dense model, maintaining local high entropy exhibits a superior scaling trend compared to the baseline. In conclusion, strategies that promote precision, either globally via generously rewarding positive tokens or locally by aggressively penalizing tail negative tokens, enables the model to converge to higher-quality solutions, potentially providing a better exploration space for RL.
To better understand the performance divergence observed during RL, we analyze the evolution of policy entropy and response length throughout the training process, as illustrated in Figure 8. Contrary to the expectation that higher entropy maintains diversity, setting a higher $\beta$ leads to rapid entropy collapse during the early stages of training. Coinciding with this collapse, the response length decreases drastically, indicating a suppression of the reasoning capability. In contrast, local high-entropy configurations exhibit greater stability. These settings effectively prevent entropy collapse, maintaining a robust policy distribution from the onset. They demonstrate a smooth and continuous activation of long reasoning capabilities, allowing for a steady increase in generation length and reasoning depth without the recovery lag observed in global high entropy settings.
Figure 8: Changes of entropy and response length during RL training across various actor models, developed based on 4B dense, 10B-A0.5B MoE, and 20B-A1B MoE architectures under different configurations.
3.6 $\text{Pass@}k$ Analysis of Base Models
Moreover, we analyze the $\text{Pass@}k$ curves as $k$ increases to estimate the upper bound of the capability of base models. This metric relies on a delicate equilibrium between solution precision and diversity. As shown in Figure 9, maximizing global diversity (high entropy) does not inherently yield higher $\text{Pass@}k$ curves. Instead, superior $\text{Pass@}k$ scores in mathematics and coding tasks are achieved by prioritizing precision. Crucially, we observe that this low-entropy setting does not lead to a collapse in output diversity. Rather, it maintains sufficient variation to cover the solution space. Furthermore, the data indicate that promoting local diversity also yields better results. This suggests that while models benefit from high precision, they simultaneously benefit from targeted local exploration.
Figure 9: $\text{Pass@}k$ curve of base models on mathematics reasoning and code generation tasks, developed based on 4B dense, 10B-A0.5B MoE and 20B-A1B MoE models under different configurations.
4 Related Works
4.1 Weighted Cross-Entropy Loss
The standard cross-entropy objective can be generalized within a policy-gradient framework, where it is equivalent to optimizing a sparse reward defined as $r_{\text{CE}}(s_{t},a_{t})=\mathbbm{1}(a_{t}=x_{t})\pi_{\theta}(a_{t}\mid s_{t})^{-1}$. Existing modifications to this objective include smooth loss (label smoothing), which encourages diversity by allocating a uniform probability mass to all positive tokens, and focal loss (Lin et al., 2018), which down-weights easy examples via $w_{t}=(1-\pi_{\theta}(x_{t}\mid s_{t}))^{\gamma}$. Our proposed generalized training objective can also formulate these established variations. In this paper, we specifically explore two different reward configurations within this framework. Firstly, we introduce a modified positive reward, which is equivalent to applying a state-dependent weight $w_{t}=\pi_{\theta}(x_{t}\mid s_{t})^{1-(1-\pi_{\theta}(x_{t}\mid s_{t}))^{\beta}}$ to the standard cross-entropy. In addition, we incorporate TopK-based negative shaping, which explicitly controls local entropy by assigning non-zero rewards to selected actions with $a_{t}\neq x_{t}$.
4.2 Next Token Reasoning
Treating each token emission as a distinct episode ensures that the reward depends only on the immediate state-action pair $(s_{t}, a_{t})$, thereby preserving unbiased credit assignment.  The framework is naturally compatible with architectures that perform iterative internal computation prior to token emission, including latent-reasoning models (Zelikman et al., 2024) and loop transformers (Dehghani et al., 2019; Zhu et al., 2025a). Although each episode terminates at token emission, the state $s_{t}$ may encode the outcome of internal refinement cycles. Our reward design can serve as an uncertainty-aware learning signal that can be combined with adaptive computation policies to allocate additional internal processing steps in uncertain contexts. By explicitly shaping positive and negative token-level rewards within a single-step policy-gradient framework, we provide a general and controllable mechanism that natively supports reasoning-oriented architectures through principled reward design.
5 Conclusion
This study establishes a theoretical bridge between next-token prediction and RL by interpreting cross-entropy loss as a specific instance of policy gradient optimization. To exploit this connection, we introduce a generalized pre-training objective that utilizes a reward-shaping strategy with positive scaling and rank-aware negative rewards. Our experiments across multiple architectures and scales reveal that regulating the diversity-precision trade-off during pre-training modulates token entropy. Our findings indicate that precision-focused strategies (e.g., global entropy reduction or tail-token suppression) yield superior scaling for the subsequent RL stage. These insights provide a novel perspective on optimizing pre-training for long CoT reasoning, suggesting new directions for sophisticated reward shaping in LLM development.
Appendix A Experiment Details for Pre-Training and Mid-Training
A.1 Implementation Details
For both the pre-training and mid-training phases, we employ the AdamW (Loshchilov & Hutter, 2017) optimizer, implementing a weight decay of 0.1 and applying gradient clipping at 1.0. Throughout these stages, we utilize a warmup-stable-decay learning rate schedule with a global batch size of 16M. During the stable pre-training stage, which encompasses 500B tokens, the learning rate warms up over 2000 steps before stabilizing at $3\times 10^{-4}$. Subsequently, we perform mid-training on an additional 100B tokens, gradually decaying the learning rate from $3\times 10^{-4}$ to $3\times 10^{-5}$. We set the maximum sequence length to 4096 during pre-training and extend it to 16384 for the mid-training stage. To support long-context modeling during mid-training, we increase the base frequency of RoPE (Su et al., 2024) from $1e^{4}$ to $1e^{6}$.
A.2 Model Architecture
Building upon the Qwen3 (Yang et al., 2025) architectures, we perform our experiments utilizing both dense and MoE architectures. Notably, we adopt an auxiliary loss free approach (Liu et al., 2024) for the training of the MoE models. Detailed architecture settings are provided in Table 1, where $E$ denotes the total number of experts and $E_{a}$ denotes the number of active experts.
Table 1: Detailed architectures settings of dense and MoE models.
Model $n_{\text{layer}}$ $d_{\text{model}}$ $d_{\text{ffn}}$ $d_{\text{expert}}$ $n_{\text{head}}$ $n_{\text{kvhead}}$ $E$ $E_{a}$
1B Dense 28 1536 4608 - 16 4 - -
4B Dense 36 2560 9728 - 32 8 - -
5B-A0.3B MoE 12 1024 - 320 32 4 384 12
10B-A0.5B MoE 16 1536 - 320 32 4 384 12
20B-A1B MoE 24 1536 - 480 32 4 384 12
A.3 Experiment Results
We report comprehensive evaluation results to demonstrate performance progression throughout the training process. Tables 2 to 11 present the pre-training results across various models and different training tokens. Similarly, Tables 12 to 17 summarize the performance metrics for the mid-training stage.
Appendix B Experiment Details for RL
B.1 Implementation Details
For RLVR on mathematical reasoning tasks, we employ the on-policy GRPO algorithm (Shao et al., 2024) without KL regularization. Following Yu et al. (2025), we incorporate clip-higher and dynamic sampling strategies to stabilize training. The process is conducted in two stages: an initial 700 steps with a sequence length of 8K, followed by continued training at a sequence length of 16K. We maintain a batch size of 128 and a constant learning rate of $1\times 10^{-6}$ for two stages. During training, we sample 16 outputs per prompt at a temperature of 1.0.
B.2 Experiment Results
We provide detailed evaluation results to illustrate performance trajectories during the RL process. Tables 18 to 32 display the RL results across different models and training steps.
Table 2: Pre-Training performance comparison across different $\beta$ based on 1B dense models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=-0.25;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0.50;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$
# Pre-Trained Tokens 125B 250B 375B 500B 125B 250B 375B 500B 125B 250B 375B 500B
General Knowledge MMLU (Acc.) 20.47 23.65 24.06 25.12 20.85 23.39 24.04 26.62 21.71 24.26 24.37 25.51
MMLU-Pro (Acc.) 8.01 9.25 9.54 10.06 7.63 8.23 7.98 9.02 7.85 8.55 9.44 9.59
NaturalQuestions (EM) 3.02 4.34 4.88 5.35 2.74 4.07 4.52 5.79 2.69 3.77 4.18 5.21
TriviaQA (EM) 8.45 12.32 14.75 16.03 8.65 12.63 13.53 16.25 8.35 12.39 14.24 16.49
Average 9.99 12.39 13.31 14.14 9.97 12.08 12.52 14.42 10.15 12.24 13.06 14.20
Commonsense Reasoning Hellaswag (Acc.) 38.24 44.17 46.17 47.49 38.96 44.42 46.26 48.33 38.88 43.95 46.83 48.06
SIQA (Acc.) 38.43 42.22 40.53 39.15 40.02 40.33 41.50 42.32 39.36 40.84 42.02 42.32
PIQA (Acc.) 67.36 69.53 69.70 71.27 67.90 69.75 71.00 71.11 67.57 69.59 70.62 70.02
WinoGrande (Acc.) 51.70 49.96 52.09 51.62 52.49 52.17 52.80 53.83 53.43 54.54 54.70 53.28
OpenBookQA (Acc.) 31.40 32.40 31.80 32.20 29.80 32.20 33.60 32.80 30.00 31.60 31.80 33.40
CommonsenseQA (Acc.) 19.66 19.00 21.13 20.80 19.57 21.79 19.82 20.07 20.15 19.41 20.72 20.56
Average 41.13 42.88 43.57 43.76 41.46 43.44 44.16 44.74 41.57 43.32 44.45 44.61
Knowledge Average 25.56 27.64 28.44 28.95 25.71 27.76 28.34 29.58 25.86 27.78 28.75 29.40
Logic Reasoning ARC-e (Acc.) 52.74 55.30 59.43 59.26 51.64 55.68 56.70 58.80 49.24 55.64 55.51 58.88
ARC-c (Acc.) 25.34 27.47 30.03 30.97 27.30 29.44 29.52 29.01 25.51 27.05 27.30 27.90
BBH (Acc.) 23.47 23.41 26.46 26.68 22.13 22.56 25.68 27.34 25.13 22.49 24.99 26.37
Average 33.85 35.39 38.64 38.97 33.69 35.89 37.30 38.38 33.29 35.06 35.93 37.72
Mathematics GSM8K (Pass@64) 40.06 43.01 46.15 49.52 41.09 48.78 48.77 48.76 43.52 46.65 46.58 49.98
MATH-500 (Pass@64) 34.16 37.09 39.79 38.15 33.41 35.43 37.10 38.49 30.41 39.64 40.80 38.97
Minerva (Pass@64) 14.62 16.19 17.07 15.76 16.44 16.71 14.80 16.43 14.94 15.32 16.54 15.99
OlympiadBench (Pass@64) 20.39 22.72 23.50 22.35 21.06 21.01 21.85 21.83 20.83 22.91 23.00 22.91
Average 27.31 29.75 31.63 31.45 28.00 30.48 30.63 31.38 27.43 31.13 31.73 31.96
Coding HumanEval+ (Pass@64) 8.06 13.03 15.46 15.81 8.71 12.15 12.72 15.95 7.68 11.92 13.67 14.55
MBPP+ (Pass@64) 21.39 33.69 40.36 41.70 20.63 34.37 37.53 41.67 16.58 31.37 38.18 39.26
Average 14.73 23.36 27.91 28.76 14.67 23.26 25.13 28.81 12.13 21.65 25.93 26.91
Reasoning Average 25.29 29.50 32.73 33.06 25.45 29.88 31.02 32.86 24.28 29.28 31.20 32.19
Average 25.40 28.76 31.01 31.41 25.56 29.03 29.95 31.55 24.91 28.68 30.22 31.08
Table 3: Pre-Training performance comparison across different $\tilde{\lambda}$ and $\hat{\lambda}$ based on 1B dense models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=-0.1$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0.1;\;\hat{\lambda}=0$
# Pre-Trained Tokens 125B 250B 375B 500B 125B 250B 375B 500B 125B 250B 375B 500B
General Knowledge MMLU (Acc.) 21.61 23.71 24.04 26.06 20.85 23.39 24.04 26.62 23.59 22.92 24.63 25.89
MMLU-Pro (Acc.) 8.34 9.32 8.42 8.67 7.63 8.23 7.98 9.02 9.04 8.69 8.50 9.23
NaturalQuestions (EM) 2.66 4.71 4.90 5.21 2.74 4.07 4.52 5.79 3.05 4.79 4.74 5.15
TriviaQA (EM) 8.61 12.70 14.95 15.59 8.65 12.63 13.53 16.25 8.76 11.18 14.08 15.50
Average 10.31 12.61 13.08 13.88 9.97 12.08 12.52 14.42 11.11 11.90 12.99 13.94
Commonsense Reasoning Hellaswag (Acc.) 38.61 44.06 46.40 48.17 38.96 44.42 46.26 48.33 38.87 43.72 46.67 48.26
SIQA (Acc.) 39.51 38.84 40.63 40.48 40.02 40.33 41.50 42.32 38.79 40.53 41.30 42.37
PIQA (Acc.) 66.70 69.64 70.84 70.95 67.90 69.75 71.00 71.11 67.30 69.48 71.60 71.11
WinoGrande (Acc.) 50.20 50.36 51.54 52.33 52.49 52.17 52.80 53.83 48.93 51.38 54.78 52.64
OpenBookQA (Acc.) 29.40 31.60 32.20 32.40 29.80 32.20 33.60 32.80 30.80 30.20 32.20 31.40
CommonsenseQA (Acc.) 19.49 19.08 20.39 18.92 19.57 21.79 19.82 20.07 20.64 19.00 18.84 23.34
Average 40.65 42.26 43.67 43.88 41.46 43.44 44.16 44.74 40.89 42.39 44.23 44.85
Knowledge Average 25.48 27.44 28.37 28.88 25.71 27.76 28.34 29.58 26.00 27.14 28.61 29.40
Logic Reasoning ARC-e (Acc.) 52.36 57.28 59.09 58.46 51.64 55.68 56.70 58.80 50.21 55.01 58.38 60.35
ARC-c (Acc.) 27.65 29.86 28.84 29.78 27.30 29.44 29.52 29.01 26.02 28.16 30.46 30.20
BBH (Acc.) 24.48 24.33 25.19 23.42 22.13 22.56 25.68 27.34 23.32 24.42 24.05 25.59
Average 34.83 37.16 37.71 37.22 33.69 35.89 37.30 38.38 33.18 35.86 37.63 38.71
Mathematics GSM8K (Pass@64) 38.32 44.09 47.21 47.38 41.09 48.78 48.77 48.76 41.85 43.30 45.25 50.04
MATH-500 (Pass@64) 34.17 36.42 38.42 39.69 33.41 35.43 37.10 38.49 33.52 38.01 37.35 37.29
Minerva (Pass@64) 14.35 16.68 16.20 17.07 16.44 16.71 14.80 16.43 15.11 16.89 17.21 16.87
OlympiadBench (Pass@64) 20.44 22.29 21.95 23.76 21.06 21.01 21.85 21.83 21.12 21.24 20.77 21.62
Average 26.82 29.87 30.95 31.98 28.00 30.48 30.63 31.38 27.90 29.86 30.15 31.46
Coding HumanEval+ (Pass@64) 8.13 13.04 15.12 14.69 8.71 12.15 12.72 15.95 8.84 12.34 16.91 16.39
MBPP+ (Pass@64) 19.08 30.08 38.37 43.02 20.63 34.37 37.53 41.67 19.25 32.66 40.90 41.62
Average 13.61 21.56 26.75 28.86 14.67 23.26 25.13 28.81 14.05 22.50 28.91 29.01
Reasoning Average 25.09 29.53 31.80 32.68 25.45 29.88 31.02 32.86 25.04 29.41 32.23 33.06
Average 25.24 28.69 30.43 31.16 25.56 29.03 29.95 31.55 25.43 28.50 30.78 31.59
Table 4: Pre-Training performance comparison across different $\beta$ based on 4B dense models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=-0.25;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0.50;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$
# Pre-Trained Tokens 125B 250B 375B 500B 125B 250B 375B 500B 125B 250B 375B 500B
General Knowledge MMLU (Acc.) 24.98 30.94 34.38 36.38 25.76 29.91 33.61 36.17 24.45 30.14 32.84 36.55
MMLU-Pro (Acc.) 8.81 11.11 11.99 13.47 9.25 10.90 11.71 12.23 8.48 9.93 10.96 12.63
NaturalQuestions (EM) 6.87 9.11 10.89 12.58 6.65 9.75 10.89 12.22 6.32 8.92 10.28 11.69
TriviaQA (EM) 17.26 25.14 30.39 33.39 17.91 25.94 27.80 33.83 17.19 24.26 30.10 32.90
Average 14.48 19.08 21.91 23.96 14.89 19.13 21.00 23.61 14.11 18.31 21.05 23.44
Commonsense Reasoning Hellaswag (Acc.) 50.36 57.81 60.80 63.01 50.22 57.00 60.50 62.91 39.15 57.49 60.98 62.87
SIQA (Acc.) 42.99 44.27 44.32 45.44 42.37 43.35 43.65 44.73 41.40 41.71 43.24 44.78
PIQA (Acc.) 71.98 74.65 75.52 75.57 72.42 74.70 75.35 76.28 71.76 74.43 74.32 75.46
WinoGrande (Acc.) 52.72 56.12 57.22 59.04 53.67 55.80 56.67 58.72 53.67 55.80 56.67 58.72
OpenBookQA (Acc.) 33.00 36.00 36.00 36.80 33.40 36.40 37.00 36.00 34.00 36.00 37.80 39.40
CommonsenseQA (Acc.) 21.13 29.48 37.43 49.63 20.39 28.26 47.91 52.91 18.43 29.32 37.10 48.98
Average 45.36 49.72 51.88 54.92 45.41 49.25 53.51 55.26 42.85 48.88 51.91 55.03
Knowledge Average 29.92 34.40 36.90 39.44 30.15 34.19 37.26 39.44 28.48 33.59 36.48 39.23
Logic Reasoning ARC-e (Acc.) 61.15 66.84 69.44 70.29 60.27 63.05 67.97 67.55 58.96 64.98 65.82 66.75
ARC-c (Acc.) 31.83 35.49 36.77 37.80 30.55 33.96 36.77 37.54 31.91 33.61 37.12 37.12
BBH (Acc.) 26.14 26.32 30.36 31.65 26.54 27.42 29.53 28.29 25.02 26.95 28.17 29.78
Average 39.71 42.88 45.52 46.58 39.12 41.48 44.76 44.46 38.63 41.85 43.70 44.55
Mathematics GSM8K (Pass@64) 49.66 62.66 67.23 71.19 48.43 60.24 68.75 71.26 50.39 61.71 69.88 71.98
MATH-500 (Pass@64) 39.07 47.73 48.62 51.14 38.19 48.15 48.88 51.54 40.86 46.72 48.23 51.67
Minerva (Pass@64) 15.48 20.73 19.37 20.07 15.09 19.22 19.68 20.63 16.33 18.45 17.64 19.85
OlympiadBench (Pass@64) 22.24 24.87 24.08 24.18 21.79 25.53 23.70 24.84 22.23 23.81 25.33 24.87
Average 31.61 39.00 39.83 41.65 30.88 38.29 40.25 42.07 32.45 37.67 40.27 42.09
Coding HumanEval+ (Pass@64) 17.11 22.94 27.79 31.29 17.32 23.22 30.08 29.13 16.73 24.36 27.72 28.52
MBPP+ (Pass@64) 44.07 59.66 65.54 65.63 40.66 56.64 65.69 66.29 43.90 59.74 65.07 65.65
Average 30.59 41.30 46.67 48.46 28.99 39.93 47.89 47.71 30.32 42.05 46.40 47.09
Reasoning Average 33.97 41.06 44.00 45.56 33.00 39.90 44.30 44.75 33.80 40.52 43.46 44.58
Average 32.35 38.40 41.16 43.11 31.86 37.61 41.48 42.62 31.67 37.75 40.66 42.44
Table 5: Pre-Training performance comparison across different $\tilde{\lambda}$ and $\hat{\lambda}$ based on 4B dense models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=-0.1$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0.1;\;\hat{\lambda}=0$
# Pre-Trained Tokens 125B 250B 375B 500B 125B 250B 375B 500B 125B 250B 375B 500B
General Knowledge MMLU (Acc.) 23.79 30.57 33.71 34.80 25.76 29.91 33.61 36.17 23.02 31.66 33.80 35.43
MMLU-Pro (Acc.) 9.57 10.37 11.50 11.83 9.25 10.90 11.71 12.23 8.44 11.37 11.87 12.20
NaturalQuestions (EM) 6.15 9.92 11.41 11.77 5.26 9.34 10.86 12.35 6.65 9.75 10.89 12.22
TriviaQA (EM) 17.19 26.53 31.13 33.59 17.91 25.94 27.80 33.83 16.93 25.99 30.57 33.29
Average 14.18 19.35 21.94 23.00 14.89 19.13 21.00 23.61 13.41 19.59 21.78 23.32
Commonsense Reasoning Hellaswag (Acc.) 50.56 57.22 60.25 62.27 50.22 57.00 60.50 62.91 50.21 58.09 61.30 62.09
SIQA (Acc.) 42.17 42.48 44.11 46.98 42.37 43.35 43.65 44.73 41.40 44.52 45.39 45.29
PIQA (Acc.) 71.27 74.37 74.16 75.19 72.42 74.70 75.35 76.28 71.16 73.88 73.56 76.12
WinoGrande (Acc.) 54.62 56.43 59.04 58.88 53.67 55.80 56.67 58.72 53.91 57.54 58.25 59.91
OpenBookQA (Acc.) 34.40 38.60 38.00 37.80 33.40 36.40 37.00 36.00 32.80 35.80 36.40 37.40
CommonsenseQA (Acc.) 19.41 30.06 39.80 46.76 20.39 28.26 47.91 52.91 20.56 30.06 41.52 46.93
Average 45.41 49.86 52.56 54.65 45.41 49.25 53.51 55.26 45.01 49.98 52.74 54.62
Knowledge Average 29.79 34.60 37.25 38.82 30.15 34.19 37.26 39.44 29.21 34.79 37.26 38.97
Logic Reasoning ARC-e (Acc.) 59.51 65.32 68.31 68.14 60.27 63.05 67.97 67.55 60.19 67.05 68.48 67.63
ARC-c (Acc.) 31.57 35.15 36.95 37.20 30.55 33.96 36.77 37.54 31.66 35.75 38.23 37.20
BBH (Acc.) 27.38 28.08 29.15 29.15 26.54 27.42 29.53 28.29 26.22 27.15 30.32 28.80
Average 39.49 42.85 44.80 44.83 39.12 41.48 44.76 44.46 39.36 43.32 45.68 44.54
Mathematics GSM8K (Pass@64) 46.40 59.61 62.93 70.92 48.43 60.24 68.75 71.26 48.09 58.17 61.84 70.94
MATH-500 (Pass@64) 38.86 47.09 48.23 52.47 38.19 48.15 48.88 51.54 38.16 43.83 46.77 50.45
Minerva (Pass@64) 16.60 17.99 17.90 21.27 15.09 19.22 19.68 20.63 16.67 18.06 19.47 20.54
OlympiadBench (Pass@64) 21.80 23.30 24.40 24.19 21.79 25.53 23.70 24.84 22.92 25.74 25.16 25.62
Average 30.92 37.00 38.37 42.21 30.88 38.29 40.25 42.07 31.46 36.45 38.31 41.89
Coding HumanEval+ (Pass@64) 16.71 23.77 27.27 29.41 17.32 23.22 30.08 29.13 15.42 20.41 29.56 31.08
MBPP+ (Pass@64) 42.24 56.78 63.88 66.01 40.66 56.64 65.69 66.29 38.19 57.32 64.74 65.65
Average 29.48 40.28 45.58 47.71 28.99 39.93 47.89 47.71 26.81 38.87 47.15 48.37
Reasoning Average 33.29 40.04 42.91 44.92 33.00 39.90 44.30 44.75 32.54 39.54 43.71 44.93
Average 31.89 37.87 40.65 42.48 31.86 37.61 41.48 42.62 29.21 37.64 41.13 42.55
Table 6: Pre-Training performance comparison across different $\beta$ based on 5B-A0.3B MoE models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=-0.25;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0.50;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$
# Pre-Trained Tokens 125B 250B 375B 500B 125B 250B 375B 500B 125B 250B 375B 500B
General Knowledge MMLU (Acc.) 22.70 27.21 29.36 29.77 23.10 25.12 29.17 31.19 24.13 26.99 28.40 30.34
MMLU-Pro (Acc.) 8.80 8.56 9.42 11.16 9.47 8.89 10.06 11.87 8.62 9.83 9.15 10.36
NaturalQuestions (EM) 5.21 7.45 8.45 10.06 5.15 7.51 8.73 10.42 5.76 7.48 8.78 10.17
TriviaQA (EM) 15.05 22.23 25.51 28.08 14.93 22.03 25.27 28.25 13.75 21.27 25.96 28.00
Average 12.94 16.36 18.19 19.77 13.16 15.89 18.31 20.43 13.07 16.39 18.07 19.72
Commonsense Reasoning Hellaswag (Acc.) 47.89 54.08 56.70 57.78 48.54 54.63 56.88 57.41 48.58 54.31 56.81 58.13
SIQA (Acc.) 40.53 41.25 42.68 43.14 41.25 42.27 43.76 42.68 40.28 42.43 42.01 43.19
PIQA (Acc.) 71.49 72.63 73.67 75.24 71.27 73.18 74.43 74.76 71.60 73.45 74.59 74.59
WinoGrande (Acc.) 52.88 53.35 57.54 57.38 50.43 53.20 56.51 56.20 52.01 54.22 55.09 56.43
OpenBookQA (Acc.) 30.80 35.60 34.00 36.60 31.20 32.80 34.40 33.60 33.80 34.00 35.80 35.60
CommonsenseQA (Acc.) 20.64 27.27 33.25 38.57 18.67 22.93 28.50 35.38 19.41 22.69 25.88 31.37
Average 44.04 47.36 49.64 51.45 43.56 46.50 49.08 50.01 44.28 46.85 48.36 49.89
Knowledge Average 28.49 31.86 33.91 35.61 28.36 31.19 33.69 35.22 28.67 31.62 33.22 34.80
Logic Reasoning ARC-e (Acc.) 57.79 60.98 62.12 62.08 58.54 62.92 64.48 63.34 58.12 64.02 63.55 63.38
ARC-c (Acc.) 27.47 32.25 32.94 34.39 30.80 34.81 35.67 35.15 28.67 33.36 34.13 34.56
BBH (Acc.) 23.30 26.80 27.12 27.85 23.97 25.79 27.71 27.03 25.94 26.69 27.00 27.49
Average 36.19 40.01 40.73 41.44 37.77 41.17 42.62 41.84 37.58 41.36 41.56 41.81
Mathematics GSM8K (Pass@64) 52.99 59.43 64.26 66.78 51.48 59.15 61.81 65.65 51.36 56.21 62.12 65.04
MATH-500 (Pass@64) 39.15 43.06 48.46 52.00 37.57 43.69 46.91 51.19 38.99 42.89 47.80 48.26
Minerva (Pass@64) 17.91 18.98 18.79 21.79 15.47 16.93 18.00 17.92 14.81 18.87 17.61 19.36
OlympiadBench (Pass@64) 22.79 23.65 24.29 25.62 23.50 24.14 23.40 24.42 23.73 23.10 25.18 26.00
Average 33.21 36.28 38.95 41.55 32.01 35.98 37.53 39.80 32.22 35.27 38.18 39.67
Coding HumanEval+ (Pass@64) 18.11 23.31 26.20 28.44 17.52 22.79 26.15 29.32 17.14 22.20 27.53 27.95
MBPP+ (Pass@64) 37.25 55.06 58.57 60.06 39.81 54.29 55.93 57.77 38.75 52.50 56.45 59.97
Average 27.68 39.19 42.39 44.25 28.67 38.54 41.04 43.55 27.95 37.35 41.99 43.96
Reasoning Average 32.36 38.49 40.69 42.41 32.81 38.56 40.40 41.73 32.58 37.99 40.58 41.81
Average 30.81 35.84 37.98 39.69 31.03 35.62 37.72 39.12 31.02 35.44 37.63 39.01
Table 7: Pre-Training performance comparison across different $\tilde{\lambda}$ and $\hat{\lambda}$ based on 5B-A0.3B MoE models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=-0.1$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0.1;\;\hat{\lambda}=0$
# Pre-Trained Tokens 125B 250B 375B 500B 125B 250B 375B 500B 125B 250B 375B 500B
General Knowledge MMLU (Acc.) 22.24 26.11 26.41 29.99 23.10 25.12 29.17 31.19 22.20 26.45 29.44 31.37
MMLU-Pro (Acc.) 7.55 8.74 8.82 10.36 9.47 8.89 10.06 11.87 8.18 9.63 10.11 11.02
NaturalQuestions (EM) 5.60 7.34 7.37 8.84 5.15 7.51 8.73 10.42 5.21 7.95 8.31 9.81
TriviaQA (EM) 14.54 21.20 25.70 27.26 14.93 22.03 25.27 28.25 15.09 21.84 25.85 28.91
Average 12.48 15.85 17.08 19.11 13.16 15.89 18.31 20.43 12.67 16.47 18.43 20.28
Commonsense Reasoning Hellaswag (Acc.) 48.44 53.64 56.72 57.58 48.54 54.63 56.88 57.41 48.51 53.89 56.42 57.33
SIQA (Acc.) 40.48 43.55 43.65 42.84 41.25 42.27 43.76 42.68 38.84 41.50 41.50 41.81
PIQA (Acc.) 69.48 72.96 73.39 74.37 71.27 73.18 74.43 74.76 71.55 74.43 74.43 74.92
WinoGrande (Acc.) 52.80 54.85 55.01 55.56 50.43 53.20 56.51 56.20 52.88 54.22 56.91 55.96
OpenBookQA (Acc.) 32.00 31.80 33.40 34.20 31.20 32.80 34.40 33.60 32.20 34.40 34.40 36.40
CommonsenseQA (Acc.) 19.74 23.67 20.15 30.30 18.67 22.93 28.50 35.38 20.07 24.65 27.76 31.86
Average 43.82 46.75 47.05 49.14 43.56 46.50 49.08 50.01 44.01 47.18 48.57 49.71
Knowledge Average 28.15 31.30 32.06 34.13 28.36 31.19 33.69 35.22 28.34 31.82 33.50 35.00
Logic Reasoning ARC-e (Acc.) 57.15 61.49 63.43 63.01 58.54 62.92 64.48 63.34 58.63 62.25 62.42 62.08
ARC-c (Acc.) 30.20 33.28 34.73 33.62 30.80 34.81 35.67 35.15 31.06 34.73 34.04 34.98
BBH (Acc.) 24.56 24.60 28.29 26.45 23.97 25.79 27.71 27.03 22.33 26.28 28.09 27.74
Average 37.30 39.79 42.15 41.03 37.77 41.17 42.62 41.84 37.34 41.09 41.52 41.60
Mathematics GSM8K (Pass@64) 50.99 59.44 62.10 64.61 51.48 59.15 61.81 65.65 48.16 58.15 60.69 59.90
MATH-500 (Pass@64) 39.91 46.65 48.92 50.65 37.57 43.69 46.91 51.19 38.71 42.31 46.49 48.60
Minerva (Pass@64) 16.36 15.98 18.30 18.83 15.47 16.93 18.00 17.92 17.31 18.01 17.78 18.02
OlympiadBench (Pass@64) 22.16 23.01 25.23 25.54 23.50 24.14 23.40 24.42 22.62 24.02 24.57 24.19
Average 32.36 36.27 38.64 39.91 32.01 35.98 37.53 39.80 31.70 35.62 37.38 37.68
Coding HumanEval+ (Pass@64) 14.01 22.49 27.53 27.00 17.52 22.79 26.15 29.32 15.27 21.36 24.62 27.88
MBPP+ (Pass@64) 38.32 52.88 57.44 60.50 39.81 54.29 55.93 57.77 40.39 49.00 56.04 60.17
Average 26.17 37.69 42.49 43.75 28.67 38.54 41.04 43.55 27.83 35.18 40.33 44.03
Reasoning Average 31.94 37.92 41.09 41.56 32.81 38.56 40.40 41.73 32.29 37.30 39.74 41.10
Average 30.43 35.27 37.48 38.59 31.03 35.62 37.72 39.12 30.71 35.11 37.25 38.66
Table 8: Pre-Training performance comparison across different $\beta$ based on 10B-A0.5B MoE models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=-0.25;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0.50;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$
# Pre-Trained Tokens 125B 250B 375B 500B 125B 250B 375B 500B 125B 250B 375B 500B
General Knowledge MMLU (Acc.) 26.13 32.74 35.09 36.60 25.35 29.14 33.50 35.12 26.05 30.71 34.03 35.48
MMLU-Pro (Acc.) 9.44 10.21 11.38 12.26 8.24 10.01 11.64 12.15 8.81 11.32 10.97 11.90
NaturalQuestions (EM) 6.79 10.36 12.47 13.99 7.06 10.47 11.41 11.96 7.84 10.58 11.36 13.77
TriviaQA (EM) 20.69 29.78 35.25 37.95 19.94 29.86 33.89 37.31 20.73 29.64 35.31 39.13
Average 15.76 20.77 23.55 25.20 15.15 19.87 22.61 24.14 15.86 20.56 22.92 25.07
Commonsense Reasoning Hellaswag (Acc.) 52.86 59.47 61.65 63.26 53.00 59.27 62.00 63.71 53.09 59.17 61.96 63.48
SIQA (Acc.) 41.61 43.45 45.14 45.60 42.02 42.73 43.45 45.09 41.91 42.84 44.63 43.76
PIQA (Acc.) 73.45 74.32 74.76 75.68 72.74 74.59 75.52 76.71 74.10 74.92 76.55 76.88
WinoGrande (Acc.) 54.06 56.59 56.27 58.64 52.88 56.12 59.04 58.88 53.83 56.04 58.80 59.04
OpenBookQA (Acc.) 34.40 35.00 35.40 37.40 32.80 35.60 35.60 38.20 33.20 37.60 37.20 39.40
CommonsenseQA (Acc.) 20.07 34.64 44.80 51.60 22.52 33.74 35.54 43.90 21.54 32.02 37.51 43.16
Average 46.08 50.58 53.00 55.36 45.99 50.34 51.86 54.42 46.28 50.43 52.78 54.29
Knowledge Average 30.92 35.68 38.28 40.28 30.57 35.11 37.23 39.28 31.07 35.50 37.85 39.68
Logic Reasoning ARC-e (Acc.) 60.94 65.91 66.84 70.29 63.01 67.09 67.63 67.00 62.37 66.37 68.77 69.32
ARC-c (Acc.) 31.91 37.20 35.67 37.46 32.68 35.24 36.52 37.29 32.51 36.09 37.20 38.82
BBH (Acc.) 27.34 28.24 29.35 29.30 25.21 25.37 28.15 29.44 24.51 27.37 27.74 28.29
Average 40.06 43.78 43.95 45.68 40.30 42.57 44.10 44.58 39.80 43.28 44.57 45.48
Mathematics GSM8K (Pass@64) 57.59 66.86 66.89 73.64 59.40 70.02 75.25 76.63 56.16 66.38 72.31 74.60
MATH-500 (Pass@64) 42.95 52.32 54.27 57.82 41.29 50.96 53.70 56.76 41.69 48.35 50.79 56.46
Minerva (Pass@64) 16.68 18.55 20.30 21.75 17.23 18.24 21.93 22.41 17.37 19.38 19.95 21.71
OlympiadBench (Pass@64) 22.26 24.79 24.54 24.58 20.53 22.27 22.77 24.31 22.04 25.70 24.86 26.52
Average 34.87 40.63 41.50 44.45 34.61 40.37 43.41 45.03 34.32 39.95 41.98 44.82
Coding HumanEval+ (Pass@64) 20.72 29.90 34.53 34.19 19.12 28.36 30.82 34.66 19.07 26.91 32.81 33.38
MBPP+ (Pass@64) 52.05 65.95 70.15 73.34 51.01 63.64 66.74 70.16 51.40 62.94 68.06 72.47
Average 36.39 47.93 52.34 53.77 35.07 46.00 48.78 52.41 35.24 44.93 50.44 52.93
Reasoning Average 37.11 44.11 45.93 47.97 36.66 42.98 45.43 47.34 36.45 42.72 45.66 47.74
Average 34.63 40.74 42.87 44.89 34.22 39.83 42.15 44.11 34.30 39.83 42.54 44.52
Table 9: Pre-Training performance comparison across different $\tilde{\lambda}$ and $\hat{\lambda}$ based on 10B-A0.5B MoE models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=-0.1$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0.1;\;\hat{\lambda}=0$
# Pre-Trained Tokens 125B 250B 375B 500B 125B 250B 375B 500B 125B 250B 375B 500B
General Knowledge MMLU (Acc.) 25.99 31.87 35.24 36.13 25.35 29.14 33.50 35.12 25.38 32.48 34.47 36.75
MMLU-Pro (Acc.) 9.40 11.10 12.17 13.19 8.24 10.01 11.64 12.15 8.96 11.19 12.64 12.34
NaturalQuestions (EM) 7.51 11.22 11.94 12.60 7.06 10.47 11.41 11.96 7.06 10.44 11.25 13.52
TriviaQA (EM) 20.55 29.22 33.57 37.30 19.94 29.86 33.89 37.31 21.12 29.48 33.93 37.61
Average 15.86 20.85 23.23 24.81 15.15 19.87 22.61 24.14 15.63 20.90 23.07 25.06
Commonsense Reasoning Hellaswag (Acc.) 52.70 59.23 63.30 63.68 53.00 59.27 62.00 63.71 53.06 59.75 63.67 63.38
SIQA (Acc.) 40.63 43.60 44.37 44.17 42.02 42.73 43.45 45.09 41.04 44.17 45.34 45.50
PIQA (Acc.) 73.23 74.92 75.90 76.22 72.74 74.59 75.52 76.71 72.42 74.70 75.52 76.50
WinoGrande (Acc.) 52.72 57.54 58.56 59.98 52.88 56.12 59.04 58.88 53.83 57.14 59.12 59.98
OpenBookQA (Acc.) 34.20 35.60 37.00 36.80 32.80 35.60 35.60 38.20 33.60 35.40 36.20 35.60
CommonsenseQA (Acc.) 21.38 37.92 44.64 48.89 22.52 33.74 35.54 43.90 19.25 34.81 41.03 45.29
Average 45.81 51.47 53.96 54.96 45.99 50.34 51.86 54.42 45.53 51.00 53.48 54.38
Knowledge Average 30.84 36.16 38.60 39.88 30.57 35.11 37.23 39.28 30.58 35.95 38.28 39.72
Logic Reasoning ARC-e (Acc.) 60.69 66.25 69.61 68.43 63.01 67.09 67.63 67.00 64.06 65.99 70.03 67.59
ARC-c (Acc.) 33.87 37.63 38.73 37.71 32.68 35.24 36.52 37.29 34.13 37.71 37.88 39.42
BBH (Acc.) 25.63 26.34 29.06 29.10 25.21 25.37 28.15 29.44 25.20 29.46 29.72 29.70
Average 40.06 43.41 45.80 45.08 40.30 42.57 44.10 44.58 41.13 44.39 45.88 45.57
Mathematics GSM8K (Pass@64) 55.13 69.42 71.80 77.63 59.40 70.02 75.25 76.63 55.89 65.29 71.44 75.11
MATH-500 (Pass@64) 43.56 49.61 54.32 56.39 41.29 50.96 53.70 56.76 42.58 49.50 52.46 54.75
Minerva (Pass@64) 17.23 19.23 20.27 22.60 17.23 18.24 21.93 22.41 16.07 19.55 21.59 20.38
OlympiadBench (Pass@64) 22.90 22.84 24.30 26.67 20.53 22.27 22.77 24.31 23.23 24.28 25.25 26.15
Average 34.71 40.28 42.67 45.82 34.61 40.37 43.41 45.03 34.44 39.66 42.69 44.10
Coding HumanEval+ (Pass@64) 21.82 29.67 32.92 35.18 19.12 28.36 30.82 34.66 19.54 25.32 30.97 32.96
MBPP+ (Pass@64) 48.31 63.41 68.98 71.01 51.01 63.64 66.74 70.16 50.24 60.62 68.22 71.02
Average 35.07 46.54 50.95 53.10 35.07 46.00 48.78 52.41 34.89 42.97 49.60 51.99
Reasoning Average 36.61 43.41 46.47 48.00 36.66 42.98 45.43 47.34 36.82 42.34 46.05 47.22
Average 34.30 40.51 43.32 44.75 34.22 39.83 42.15 44.11 34.33 39.78 42.94 44.22
Table 10: Pre-Training performance comparison across different $\beta$ based on 20B-A1B MoE models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=-0.25;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0.50;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$
# Pre-Trained Tokens 125B 250B 375B 500B 125B 250B 375B 500B 125B 250B 375B 500B
General Knowledge MMLU (Acc.) 28.19 35.81 39.35 41.67 29.77 35.09 39.40 41.53 27.33 35.44 38.75 40.78
MMLU-Pro (Acc.) 9.11 11.98 13.45 16.70 10.41 12.15 14.46 15.32 9.40 13.21 13.64 14.94
NaturalQuestions (EM) 8.83 13.57 15.56 18.39 10.00 14.60 16.18 17.15 9.78 14.32 16.01 18.61
TriviaQA (EM) 26.81 37.23 43.54 47.66 27.31 38.02 44.07 46.75 27.37 38.54 43.38 46.17
Average 18.24 24.65 27.98 31.11 19.37 24.97 28.53 30.19 18.47 25.38 27.95 30.13
Commonsense Reasoning Hellaswag (Acc.) 58.17 64.88 67.68 69.05 58.52 65.31 67.88 69.29 58.35 65.43 67.98 69.47
SIQA (Acc.) 42.89 45.75 45.39 47.80 42.07 44.98 46.72 46.16 42.32 44.98 45.70 46.93
PIQA (Acc.) 75.46 77.48 77.58 78.35 74.48 76.77 77.69 77.86 74.59 76.44 77.26 78.18
WinoGrande (Acc.) 55.25 57.85 60.85 62.90 55.09 60.14 59.91 61.19 55.17 57.77 62.35 64.09
OpenBookQA (Acc.) 36.80 37.80 41.00 40.40 37.00 40.20 40.40 41.20 33.20 38.80 42.40 40.20
CommonsenseQA (Acc.) 27.68 50.29 53.81 61.34 31.20 54.22 58.07 58.23 24.73 40.54 54.95 56.84
Average 49.38 55.68 57.72 59.97 49.73 56.94 58.45 58.99 48.06 53.99 58.44 59.28
Knowledge Average 33.81 40.16 42.85 45.54 34.55 40.95 43.49 44.59 33.27 39.69 43.19 44.71
Logic Reasoning ARC-e (Acc.) 65.49 70.03 72.85 72.60 62.84 70.24 70.58 72.98 66.12 69.32 72.39 74.71
ARC-c (Acc.) 34.64 39.93 41.47 43.34 34.04 40.70 44.02 43.00 35.75 40.87 41.64 43.43
BBH (Acc.) 26.69 29.10 31.84 32.50 26.88 30.16 32.07 33.82 26.36 28.81 29.93 30.75
Average 42.27 46.35 48.72 49.48 41.25 47.03 48.89 49.93 42.74 46.33 47.99 49.63
Mathematics GSM8K (Pass@64) 66.43 79.76 83.15 85.75 59.06 76.57 82.05 81.99 63.46 76.58 78.65 84.61
MATH-500 (Pass@64) 47.59 56.75 62.34 63.64 48.07 55.95 61.74 63.97 44.34 54.57 60.00 64.91
Minerva (Pass@64) 17.75 23.32 23.68 25.16 18.77 22.33 22.85 25.82 17.05 19.88 23.71 25.40
OlympiadBench (Pass@64) 25.22 25.88 28.94 27.65 26.30 26.15 28.04 28.86 24.03 25.05 26.47 27.16
Average 39.25 46.43 49.53 50.55 38.05 45.25 48.67 50.16 37.22 44.02 47.21 50.52
Coding HumanEval+ (Pass@64) 26.66 35.66 39.29 40.89 26.66 35.66 39.29 40.89 26.02 32.82 43.16 41.45
MBPP+ (Pass@64) 56.54 71.21 76.10 79.56 60.90 73.75 77.36 78.47 57.42 72.06 76.92 79.34
Average 41.60 53.44 57.70 60.23 43.44 54.60 58.64 59.74 41.72 52.44 60.04 60.40
Reasoning Average 41.04 48.74 51.98 53.42 40.91 48.96 52.07 53.28 40.56 47.60 51.75 53.51
Average 38.15 45.31 48.33 50.27 38.37 45.76 48.63 49.80 37.64 44.43 48.32 49.99
Table 11: Pre-Training performance comparison across different $\tilde{\lambda}$ and $\hat{\lambda}$ based on 20B-A1B MoE models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=-0.1$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0.1;\;\hat{\lambda}=0$
# Pre-Trained Tokens 125B 250B 375B 500B 125B 250B 375B 500B 125B 250B 375B 500B
General Knowledge MMLU (Acc.) 28.84 36.46 39.82 42.04 29.77 35.09 39.40 41.53 25.92 35.71 39.21 41.47
MMLU-Pro (Acc.) 9.77 12.92 15.24 16.20 10.41 12.15 14.46 15.32 9.53 12.26 13.68 15.35
NaturalQuestions (EM) 9.83 14.40 16.18 18.39 10.00 14.60 16.18 17.15 9.06 14.46 15.65 18.95
TriviaQA (EM) 27.44 37.48 42.76 46.19 27.31 38.02 44.07 46.75 26.42 38.38 43.95 47.16
Average 18.97 25.32 28.50 30.71 19.37 24.97 28.53 30.19 17.73 25.20 28.12 30.73
Commonsense Reasoning Hellaswag (Acc.) 57.75 64.03 67.51 69.16 58.18 64.73 68.13 69.10 58.52 65.31 67.88 69.29
SIQA (Acc.) 42.67 46.57 48.62 47.49 42.07 44.98 46.72 46.16 44.22 45.09 47.39 47.70
PIQA (Acc.) 74.86 77.42 77.31 78.24 75.41 77.20 77.09 78.73 74.48 76.77 77.69 77.86
WinoGrande (Acc.) 55.09 58.88 60.06 61.01 55.09 60.14 59.91 61.19 55.41 59.91 61.72 62.59
OpenBookQA (Acc.) 34.80 36.40 41.00 38.60 37.00 40.20 40.40 41.20 36.20 40.00 41.40 42.00
CommonsenseQA (Acc.) 30.22 50.04 56.84 59.71 31.20 54.22 58.07 58.23 22.93 44.55 53.64 58.31
Average 49.23 55.56 58.56 59.04 49.73 56.94 58.45 58.99 48.73 55.25 58.23 59.74
Knowledge Average 34.10 40.44 43.53 44.87 34.55 40.95 43.49 44.59 33.23 40.22 43.18 45.24
Logic Reasoning ARC-e (Acc.) 65.03 69.28 71.72 73.06 62.84 70.24 70.58 72.98 65.82 70.83 70.37 72.22
ARC-c (Acc.) 34.47 39.08 41.47 42.83 34.04 40.70 44.02 43.00 35.84 41.89 41.30 43.09
BBH (Acc.) 26.25 29.15 30.78 33.56 26.88 30.16 32.07 33.82 23.78 29.17 32.07 33.41
Average 41.92 45.84 47.99 49.82 41.25 47.03 48.89 49.93 41.81 47.30 47.91 49.57
Mathematics GSM8K (Pass@64) 61.97 76.10 80.08 84.39 59.06 76.57 82.05 81.99 64.57 76.15 80.86 82.14
MATH-500 (Pass@64) 46.60 56.29 63.10 65.07 48.07 55.95 61.74 63.97 47.49 56.82 61.07 65.56
Minerva (Pass@64) 19.71 20.51 22.84 23.54 18.77 22.33 22.85 25.82 17.63 22.25 23.89 23.90
OlympiadBench (Pass@64) 22.31 24.67 26.50 29.00 26.30 26.15 28.04 28.86 25.74 26.91 26.86 28.54
Average 37.65 44.39 48.13 50.50 38.05 45.25 48.67 50.16 38.86 45.53 48.17 50.04
Coding HumanEval+ (Pass@64) 24.73 34.46 41.83 40.47 25.98 35.44 39.91 41.00 25.83 36.66 40.12 40.79
MBPP+ (Pass@64) 59.21 72.34 76.80 80.47 60.90 73.75 77.36 78.47 58.90 70.87 75.22 77.69
Average 41.97 53.40 59.32 60.47 43.44 54.60 58.64 59.74 42.37 53.77 57.67 59.24
Reasoning Average 40.51 47.88 51.81 53.60 40.91 48.96 52.07 53.28 41.01 48.86 51.25 52.95
Average 37.95 44.90 48.50 50.11 38.37 45.76 48.63 49.80 37.90 45.41 48.02 49.86
Table 12: Mid-Training performance comparison across different $\beta$ based on 4B dense models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=-0.25;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0.50;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$
# Mid-Trained Tokens 25B 50B 75B 100B 25B 50B 75B 100B 25B 50B 75B 100B
General Knowledge MMLU (Acc.) 37.99 39.51 39.69 39.92 37.27 39.34 39.23 40.47 36.53 38.45 39.00 39.48
MMLU-Pro (Acc.) 16.51 18.44 19.69 19.70 15.50 17.27 17.77 18.68 14.97 16.44 17.99 18.53
NaturalQuestions (EM) 11.69 11.75 12.16 12.33 12.16 11.91 12.33 12.55 11.75 11.14 12.08 11.88
TriviaQA (EM) 32.60 32.75 34.05 34.50 31.90 32.54 33.50 34.17 32.55 32.78 34.09 34.67
Average 24.70 25.61 26.40 26.61 24.21 25.27 25.71 26.47 23.95 24.70 25.79 26.14
Commonsense Reasoning Hellaswag (Acc.) 61.59 61.70 61.95 62.41 61.25 61.72 62.14 62.09 61.68 62.01 62.54 62.51
SIQA (Acc.) 44.78 43.71 44.42 44.78 45.44 45.29 45.75 45.85 44.32 44.78 44.83 44.88
PIQA (Acc.) 75.24 76.01 75.19 75.84 75.24 75.24 75.24 75.73 74.54 74.92 74.81 74.92
WinoGrande (Acc.) 60.69 59.67 60.22 60.62 58.96 59.98 60.46 60.85 58.88 59.35 59.51 59.67
OpenBookQA (Acc.) 35.60 38.20 37.20 37.40 36.60 37.40 38.00 37.40 39.20 40.40 40.60 40.00
CommonsenseQA (Acc.) 52.09 52.42 54.38 55.77 54.13 52.99 53.97 54.87 49.14 50.04 52.58 53.07
Average 55.00 55.29 55.56 56.14 55.27 55.44 55.93 56.13 54.63 55.25 55.81 55.84
Knowledge Average 39.85 40.45 40.98 41.37 39.74 40.35 40.82 41.30 39.29 39.98 40.80 40.99
Logic Reasoning ARC-e (Acc.) 66.84 68.94 70.45 69.82 68.81 68.94 69.82 69.99 66.84 69.19 69.91 71.17
ARC-c (Acc.) 39.33 41.04 40.78 41.89 40.61 41.13 41.64 41.89 38.14 41.38 40.87 41.72
BBH (Acc.) 33.90 38.58 39.66 39.83 33.47 36.95 36.45 37.28 31.90 34.88 36.11 35.92
Average 46.69 49.52 50.30 50.51 47.63 49.01 49.30 49.72 45.63 48.48 48.96 49.60
Mathematics GSM8K (Pass@64) 84.66 88.94 91.17 91.25 85.73 89.60 92.23 92.70 83.30 88.82 91.75 92.08
MATH-500 (Pass@64) 63.69 66.70 70.73 70.41 62.42 66.14 69.48 71.35 62.42 65.49 68.17 68.97
Minerva (Pass@64) 21.88 23.67 26.82 25.99 24.70 24.38 28.22 26.72 22.86 23.90 23.95 24.99
OlympiadBench (Pass@64) 28.96 30.15 32.15 32.21 28.44 33.15 32.96 32.65 29.21 29.80 31.51 32.85
Average 49.80 52.37 55.22 54.97 50.32 53.32 55.72 55.86 49.45 52.00 53.85 54.72
Coding HumanEval+ (Pass@64) 52.37 57.74 64.85 64.33 51.64 60.52 64.95 65.89 48.07 60.60 63.31 66.30
MBPP+ (Pass@64) 79.91 82.27 85.48 86.79 77.50 82.32 84.81 85.28 77.83 82.51 83.77 83.41
Average 66.74 70.01 75.17 75.56 64.57 71.42 74.88 75.59 62.95 71.56 73.54 74.86
Reasoning Average 54.21 57.30 60.23 60.35 54.17 57.91 59.97 60.39 52.67 57.35 58.78 59.73
Average 48.46 50.56 52.53 52.76 48.40 50.89 52.31 52.75 47.32 50.40 51.59 52.23
Table 13: Mid-Training performance comparison across different $\tilde{\lambda}$ and $\hat{\lambda}$ based on 4B dense models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=-0.1$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0.1;\;\hat{\lambda}=0$
# Mid-Trained Tokens 25B 50B 75B 100B 25B 50B 75B 100B 25B 50B 75B 100B
General Knowledge MMLU (Acc.) 37.33 38.73 38.83 39.13 37.27 39.34 39.23 40.47 37.15 38.06 38.76 39.43
MMLU-Pro (Acc.) 14.17 16.61 17.57 18.23 15.50 17.27 17.77 18.68 14.17 15.58 16.41 17.50
NaturalQuestions (EM) 11.25 11.99 12.30 12.66 12.16 11.91 12.33 12.55 11.14 12.49 12.58 12.74
TriviaQA (EM) 33.29 33.89 34.80 35.32 31.90 32.54 33.50 34.17 32.73 33.50 34.17 34.59
Average 24.01 25.31 25.88 26.34 24.21 25.27 25.71 26.47 23.80 24.91 25.48 26.07
Commonsense Reasoning Hellaswag (Acc.) 61.11 61.25 61.76 62.29 61.25 61.72 62.14 62.09 61.59 61.46 61.98 62.16
SIQA (Acc.) 45.09 44.37 44.06 45.24 45.44 45.29 45.75 45.85 46.21 46.21 46.37 46.62
PIQA (Acc.) 75.46 76.01 76.22 76.12 75.24 75.24 75.24 75.73 74.59 74.76 75.41 75.41
WinoGrande (Acc.) 60.22 60.22 61.48 60.46 58.96 59.98 60.46 60.85 60.85 60.14 59.51 59.98
OpenBookQA (Acc.) 39.80 38.20 38.20 38.00 36.60 37.40 38.00 37.40 39.80 38.60 39.60 40.20
CommonsenseQA (Acc.) 46.93 51.84 53.73 53.81 54.13 52.99 53.97 54.87 45.86 47.67 49.80 50.20
Average 54.77 55.32 55.91 55.99 55.27 55.44 55.93 56.13 54.82 54.81 55.45 55.76
Knowledge Average 39.39 40.31 40.89 41.16 39.74 40.35 40.82 41.30 39.31 39.86 40.46 40.91
Logic Reasoning ARC-e (Acc.) 68.52 68.56 70.75 71.34 68.81 68.94 69.82 69.99 70.33 70.29 71.04 71.25
ARC-c (Acc.) 39.93 40.96 41.81 42.66 40.61 41.13 41.64 41.89 41.30 42.15 42.24 43.34
BBH (Acc.) 32.78 37.15 38.49 39.30 33.47 36.95 36.45 37.28 33.62 35.40 36.58 37.74
Average 47.08 48.89 50.35 51.10 47.63 49.01 49.30 49.72 48.42 49.28 49.95 50.78
Mathematics GSM8K (Pass@64) 84.38 89.90 90.43 91.69 85.73 89.60 92.23 92.70 85.99 87.51 90.68 90.11
MATH-500 (Pass@64) 64.40 70.24 72.50 72.78 62.42 66.14 69.48 71.35 63.31 67.35 70.36 70.31
Minerva (Pass@64) 21.64 23.81 25.65 26.86 24.70 24.38 28.22 26.72 23.28 22.60 25.53 25.43
OlympiadBench (Pass@64) 30.09 32.29 33.52 34.13 28.44 33.15 32.96 32.65 26.32 28.90 30.55 34.11
Average 50.13 54.06 55.53 56.37 50.32 53.32 55.72 55.86 49.73 51.59 54.28 54.99
Coding HumanEval+ (Pass@64) 50.62 58.51 64.58 65.23 51.64 60.52 64.95 65.89 51.68 58.17 63.94 65.31
MBPP+ (Pass@64) 78.24 83.29 85.36 85.22 77.50 82.32 84.81 85.28 79.31 81.59 85.17 85.23
Average 64.43 70.90 74.97 75.23 64.57 71.42 74.88 75.59 65.50 69.88 74.56 75.27
Reasoning Average 53.88 57.95 60.28 60.90 54.17 57.91 59.97 60.39 54.55 56.92 59.60 60.35
Average 48.08 50.89 52.53 53.00 48.40 50.89 52.31 52.75 48.45 50.09 51.94 52.57
Table 14: Mid-Training performance comparison across different $\beta$ based on 10B-A0.5B MoE models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=-0.25;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0.50;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$
# Mid-Trained Tokens 25B 50B 75B 100B 25B 50B 75B 100B 25B 50B 75B 100B
General Knowledge MMLU (Acc.) 36.39 37.23 37.75 38.11 36.15 37.27 37.54 37.82 35.04 36.49 36.45 37.10
MMLU-Pro (Acc.) 12.31 12.88 13.69 13.91 12.98 13.91 14.41 14.71 12.03 12.82 13.32 13.46
NaturalQuestions (EM) 12.80 13.77 14.18 14.35 11.30 12.11 12.44 12.85 11.30 12.11 12.44 12.85
TriviaQA (EM) 36.61 37.37 38.29 38.76 35.84 37.15 37.81 38.38 37.57 38.84 39.62 40.21
Average 24.53 25.31 25.98 26.28 24.07 25.11 25.55 25.94 24.40 25.41 25.87 26.18
Commonsense Reasoning Hellaswag (Acc.) 62.25 62.39 62.94 63.25 62.47 62.82 63.50 63.74 62.51 62.67 63.29 63.43
SIQA (Acc.) 43.86 44.63 44.68 45.19 45.55 43.04 43.70 44.22 44.63 44.01 44.73 44.01
PIQA (Acc.) 75.46 74.92 75.57 75.95 76.06 75.63 75.73 76.55 75.90 76.06 76.50 76.82
WinoGrande (Acc.) 58.80 59.98 60.54 60.46 58.96 58.88 59.59 59.83 58.64 59.51 60.77 60.30
OpenBookQA (Acc.) 37.40 36.80 37.60 38.00 36.60 37.20 37.00 37.20 40.00 40.40 40.20 40.60
CommonsenseQA (Acc.) 50.53 50.45 50.61 50.61 43.90 46.52 49.06 49.06 40.46 42.75 43.73 43.90
Average 54.72 54.86 55.32 55.58 53.92 54.02 54.76 55.10 53.69 54.23 54.87 54.84
Knowledge Average 39.62 40.09 40.65 40.93 39.00 39.56 40.16 40.52 39.04 39.82 40.37 40.51
Logic Reasoning ARC-e (Acc.) 68.35 68.90 69.40 70.12 68.35 68.90 69.40 70.12 68.52 70.03 69.87 70.24
ARC-c (Acc.) 38.65 40.10 41.30 40.53 39.59 40.44 42.49 42.83 39.93 40.87 41.64 41.98
BBH (Acc.) 30.70 31.64 31.84 32.05 30.64 33.68 33.59 34.14 30.69 32.42 33.25 33.44
Average 45.90 46.88 47.51 47.57 46.40 48.30 48.91 49.27 46.38 47.77 48.25 48.55
Mathematics GSM8K (Pass@64) 85.06 87.67 90.35 90.53 87.16 90.68 90.83 92.03 84.62 88.39 89.84 90.30
MATH-500 (Pass@64) 65.20 68.85 70.19 71.73 64.39 68.35 70.77 70.97 64.27 68.59 67.46 70.00
Minerva (Pass@64) 25.33 25.02 26.26 26.53 24.51 23.83 24.77 27.03 23.51 26.76 27.11 25.90
OlympiadBench (Pass@64) 29.73 30.76 32.84 33.21 28.91 30.05 33.11 32.63 28.57 29.86 32.82 33.39
Average 51.33 53.08 54.91 55.50 51.24 53.23 54.87 55.67 50.24 53.40 54.31 54.90
Coding HumanEval+ (Pass@64) 48.79 55.88 56.24 58.05 48.00 53.35 56.98 55.18 49.33 53.14 55.06 56.21
MBPP+ (Pass@64) 77.31 80.74 84.55 83.41 76.28 80.29 81.82 81.38 76.03 78.12 80.15 80.80
Average 63.05 68.31 70.40 70.73 62.14 66.82 69.40 68.28 62.68 65.63 67.61 68.51
Reasoning Average 53.43 56.09 57.61 57.93 53.26 56.12 57.73 57.74 53.10 55.60 56.72 57.32
Average 47.90 49.69 50.82 51.13 47.56 49.50 50.70 50.85 47.48 49.29 50.18 50.60
Table 15: Mid-Training performance comparison across different $\tilde{\lambda}$ and $\hat{\lambda}$ based on 10B-A0.5B MoE models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=-0.1$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0.1;\;\hat{\lambda}=0$
# Mid-Trained Tokens 25B 50B 75B 100B 25B 50B 75B 100B 25B 50B 75B 100B
General Knowledge MMLU (Acc.) 36.86 38.22 38.98 39.07 36.15 37.27 37.54 37.82 36.90 37.96 38.29 39.30
MMLU-Pro (Acc.) 14.10 15.20 15.93 16.39 12.98 13.91 14.41 14.71 13.02 15.05 14.67 14.72
NaturalQuestions (EM) 12.35 13.05 13.63 13.60 11.30 12.11 12.44 12.85 12.52 12.63 13.30 13.49
TriviaQA (EM) 35.51 37.14 37.49 38.25 35.84 37.15 37.81 38.38 36.25 36.98 38.01 38.44
Average 24.71 25.90 26.51 26.83 24.07 25.11 25.55 25.94 24.67 25.66 26.07 26.49
Commonsense Reasoning Hellaswag (Acc.) 62.67 62.96 63.12 63.50 62.47 62.82 63.50 63.74 62.93 63.07 63.42 63.73
SIQA (Acc.) 43.86 43.96 43.71 43.71 45.55 43.04 43.70 44.22 45.34 45.50 45.91 45.45
PIQA (Acc.) 76.01 75.41 75.95 76.22 76.06 75.63 75.73 76.55 76.01 75.73 76.17 76.12
WinoGrande (Acc.) 60.06 61.09 60.77 58.88 60.06 61.09 60.77 58.88 58.56 59.59 60.46 61.01
OpenBookQA (Acc.) 37.60 35.80 36.00 35.40 36.60 37.20 37.00 37.20 37.20 35.80 37.00 37.80
CommonsenseQA (Acc.) 45.62 48.89 45.86 48.48 43.90 46.52 49.06 49.06 45.45 47.83 47.91 49.63
Average 54.30 54.69 54.24 54.37 53.92 54.02 54.76 55.10 54.25 54.59 55.15 55.62
Knowledge Average 39.50 40.29 40.37 40.60 39.00 39.56 40.16 40.52 39.46 40.12 40.61 41.06
Logic Reasoning ARC-e (Acc.) 69.78 69.78 70.58 71.09 68.98 70.79 70.66 70.83 68.48 69.95 70.56 71.13
ARC-c (Acc.) 40.10 41.21 41.81 41.30 39.59 40.44 42.49 42.83 41.64 42.41 41.81 42.06
BBH (Acc.) 31.18 33.21 33.71 34.05 30.64 33.68 33.59 34.14 30.70 33.19 32.64 32.73
Average 47.02 48.07 48.70 48.81 46.40 48.30 48.91 49.27 46.94 48.52 48.34 48.64
Mathematics GSM8K (Pass@64) 87.23 90.34 91.10 91.45 87.16 90.68 90.83 92.03 84.20 87.47 89.89 90.41
MATH-500 (Pass@64) 64.74 69.10 71.01 70.77 64.39 68.35 70.77 70.97 63.07 65.99 69.52 69.96
Minerva (Pass@64) 24.50 25.23 25.50 26.67 24.51 23.83 24.77 27.03 23.05 26.98 27.02 27.41
OlympiadBench (Pass@64) 28.84 30.41 31.34 31.20 28.91 30.05 33.11 32.63 31.11 32.09 32.57 32.83
Average 51.33 53.77 54.74 55.02 51.24 53.23 54.87 55.67 50.36 53.13 54.75 55.15
Coding HumanEval+ (Pass@64) 47.08 50.89 52.11 55.14 48.00 53.35 56.98 55.18 47.36 51.56 55.66 55.62
MBPP+ (Pass@64) 78.67 79.80 81.25 84.00 76.28 80.29 81.82 81.38 76.24 80.35 82.42 83.16
Average 62.88 65.35 66.68 69.57 62.14 66.82 69.40 68.28 61.80 65.96 69.04 69.39
Reasoning Average 53.74 55.73 56.71 57.80 53.26 56.12 57.73 57.74 53.03 55.87 57.38 57.73
Average 48.05 49.55 50.17 50.92 47.56 49.50 50.70 50.85 47.60 49.57 50.67 51.06
Table 16: Mid-Training performance comparison across different $\beta$ based on 20B-A1B MoE models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=-0.25;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0.50;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$
# Mid-Trained Tokens 25B 50B 75B 100B 25B 50B 75B 100B 25B 50B 75B 100B
General Knowledge MMLU (Acc.) 43.19 44.69 45.71 46.25 43.40 45.56 46.02 46.55 41.97 43.58 44.97 45.42
MMLU-Pro (Acc.) 19.97 23.11 25.02 25.66 21.63 23.73 25.38 25.64 19.34 22.32 24.54 24.48
NaturalQuestions (EM) 16.62 18.06 17.89 18.61 17.09 17.59 17.81 17.92 16.90 16.81 17.42 17.89
TriviaQA (EM) 45.35 46.57 47.31 47.98 45.95 46.65 47.89 48.31 44.76 45.73 47.02 47.10
Average 31.28 33.11 33.98 34.63 32.02 33.38 34.28 34.60 30.74 32.11 33.49 33.72
Commonsense Reasoning Hellaswag (Acc.) 67.88 67.61 68.01 68.61 67.92 68.10 68.45 68.49 67.87 68.12 68.33 68.50
SIQA (Acc.) 44.98 44.68 45.24 44.73 45.70 46.93 46.93 46.88 44.27 46.42 45.50 45.39
PIQA (Acc.) 77.31 76.88 77.80 77.86 77.53 77.31 78.89 78.56 77.80 77.26 77.37 76.77
WinoGrande (Acc.) 61.80 61.56 62.59 63.22 62.19 62.98 63.85 62.59 62.90 64.09 62.90 63.06
OpenBookQA (Acc.) 39.40 39.40 39.60 41.00 41.20 40.80 40.80 40.20 41.20 41.20 40.80 41.00
CommonsenseQA (Acc.) 58.48 59.30 60.52 60.69 61.43 62.33 62.24 61.75 56.67 58.15 59.79 59.87
Average 58.31 58.24 58.96 59.35 59.33 59.74 60.19 59.75 58.45 59.21 59.11 59.10
Knowledge Average 44.80 45.67 46.47 46.99 45.67 46.56 47.23 47.17 44.60 45.66 46.30 46.41
Logic Reasoning ARC-e (Acc.) 70.66 72.22 72.18 73.61 73.02 74.07 74.28 74.12 74.54 74.33 74.20 74.45
ARC-c (Acc.) 43.00 43.86 45.14 45.22 45.48 47.44 47.95 47.35 45.73 46.16 46.93 46.59
BBH (Acc.) 39.99 46.95 49.15 50.25 42.34 46.74 49.90 51.33 39.30 45.65 48.58 49.59
Average 51.22 54.34 55.49 56.36 53.61 56.08 57.38 57.60 53.19 55.38 56.57 56.88
Mathematics GSM8K (Pass@64) 94.64 95.94 95.91 96.43 93.59 95.65 95.69 96.42 94.16 96.04 96.18 96.33
MATH-500 (Pass@64) 77.55 82.14 85.30 85.91 74.76 78.96 80.39 80.26 73.65 78.51 81.04 80.83
Minerva (Pass@64) 31.49 35.35 36.91 37.21 28.43 33.64 35.36 34.61 29.78 32.07 31.38 30.87
OlympiadBench (Pass@64) 38.47 47.15 46.64 47.77 35.99 38.39 38.29 40.91 38.47 45.38 52.48 51.22
Average 60.54 65.15 66.19 66.83 58.19 61.66 62.44 63.05 59.01 63.00 65.27 64.81
Coding HumanEval+ (Pass@64) 67.82 73.38 77.23 79.25 67.09 71.95 76.05 74.99 68.19 71.46 73.26 74.48
MBPP+ (Pass@64) 85.98 88.54 89.77 90.03 87.34 88.88 89.88 90.30 87.75 89.33 89.71 90.32
Average 76.90 80.96 83.50 84.64 77.22 80.42 82.96 82.65 77.97 80.39 81.49 82.40
Reasoning Average 62.89 66.82 68.39 69.28 63.01 66.05 67.59 67.76 63.39 66.26 67.78 68.03
Average 55.65 58.36 59.62 60.36 56.07 58.26 59.45 59.53 55.87 58.02 59.19 59.38
Table 17: Mid-Training performance comparison across different $\tilde{\lambda}$ and $\hat{\lambda}$ based on 20B-A1B MoE models. The highest scores at the final checkpoint across the different configurations are shown in bold.
Hyperparameters $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=-0.1$ $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$ $\beta=0;\;\tilde{\lambda}=0.1;\;\hat{\lambda}=0$
# Mid-Trained Tokens 25B 50B 75B 100B 25B 50B 75B 100B 25B 50B 75B 100B
General Knowledge MMLU (Acc.) 42.79 44.69 45.86 46.77 43.40 45.56 46.02 46.55 42.56 44.47 46.10 45.61
MMLU-Pro (Acc.) 22.57 24.29 25.91 25.83 21.63 23.73 25.38 25.64 21.57 23.55 26.33 26.15
NaturalQuestions (EM) 16.51 16.54 17.73 18.01 17.09 17.59 17.81 17.92 16.98 18.14 18.53 18.92
TriviaQA (EM) 45.43 45.85 47.29 47.82 45.95 46.65 47.89 48.31 45.97 46.90 47.72 48.35
Average 31.82 32.84 34.20 34.61 32.02 33.38 34.28 34.60 31.77 33.27 34.67 34.76
Commonsense Reasoning Hellaswag (Acc.) 67.67 68.00 68.43 68.46 67.92 68.10 68.45 68.49 68.12 67.73 68.49 68.55
SIQA (Acc.) 44.98 46.01 45.70 46.16 45.70 46.93 46.93 46.88 46.98 47.49 47.70 47.39
PIQA (Acc.) 76.88 77.86 77.86 77.80 77.53 77.31 78.89 78.56 77.37 78.45 78.51 78.73
WinoGrande (Acc.) 62.35 62.51 62.59 61.09 62.19 62.98 63.85 62.59 62.67 62.04 61.25 61.40
OpenBookQA (Acc.) 38.60 39.20 40.00 39.80 41.20 40.80 40.80 40.20 41.60 41.00 42.40 43.00
CommonsenseQA (Acc.) 59.46 59.30 60.52 60.77 61.43 62.33 62.24 61.75 57.08 59.13 58.89 59.95
Average 58.32 58.81 59.18 59.01 59.33 59.74 60.19 59.75 58.97 59.31 59.54 59.84
Knowledge Average 45.07 45.83 46.69 46.81 45.67 46.56 47.23 47.17 45.37 46.29 47.10 47.30
Logic Reasoning ARC-e (Acc.) 73.32 73.78 74.37 74.87 73.02 74.07 74.28 74.12 72.77 74.37 75.34 74.96
ARC-c (Acc.) 45.56 45.65 47.78 47.18 45.48 47.44 47.95 47.35 45.39 46.76 48.29 47.44
BBH (Acc.) 39.46 45.83 48.32 50.24 42.34 46.74 49.90 51.33 42.22 46.57 48.40 49.95
Average 52.78 55.09 56.82 57.43 53.61 56.08 57.38 57.60 53.46 55.90 57.34 57.45
Mathematics GSM8K (Pass@64) 94.18 95.34 96.39 96.51 93.59 95.65 95.69 96.42 93.81 95.16 95.47 95.76
MATH-500 (Pass@64) 76.68 78.17 80.38 80.54 74.76 78.96 80.39 80.26 76.81 78.68 83.05 84.19
Minerva (Pass@64) 29.60 31.41 33.18 37.06 28.43 33.64 35.36 34.61 32.35 34.18 39.24 37.26
OlympiadBench (Pass@64) 37.63 44.08 45.41 48.62 35.99 38.39 38.29 40.91 36.29 46.41 50.51 51.08
Average 59.52 62.25 63.84 65.68 58.19 61.66 62.44 63.05 59.82 63.61 67.07 67.07
Coding HumanEval+ (Pass@64) 66.36 70.07 71.99 71.88 67.09 71.95 76.05 74.99 63.55 69.00 74.08 75.41
MBPP+ (Pass@64) 86.64 88.79 89.46 89.97 87.34 88.88 89.88 90.30 85.21 88.59 89.32 89.22
Average 76.50 79.43 80.73 80.92 77.22 80.42 82.96 82.65 74.38 78.80 81.70 82.32
Reasoning Average 62.93 65.59 67.13 68.01 63.01 66.05 67.59 67.76 62.55 66.10 68.70 68.95
Average 55.79 57.68 58.96 59.53 56.07 58.26 59.45 59.53 55.68 58.18 60.06 60.29
Table 18: RL performance of the 4B Dense Model with $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.89 0.81 1.14 1.04 1.12 1.54 1.33 1.69 1.64 1.62
Cons@128 0.00 0.00 0.00 0.00 3.33 3.33 3.33 6.67 3.33 3.33
Pass@64 15.34 13.77 21.27 14.80 11.68 22.86 19.81 19.68 14.78 18.32
AIME25 Avg@128 0.57 0.86 0.86 1.46 1.59 2.29 1.30 2.06 2.21 2.24
Cons@128 0.00 0.00 0.00 3.33 0.00 6.67 3.33 3.33 3.33 3.33
Pass@64 14.90 11.67 10.84 10.00 16.16 10.84 12.93 14.35 10.84 14.59
AMC23 Avg@128 18.46 23.07 24.94 27.87 28.96 28.73 27.17 28.18 29.16 29.51
Cons@128 37.50 35.00 40.00 47.50 45.00 45.00 45.00 45.00 47.50 45.00
Pass@64 71.39 76.75 79.44 75.70 74.24 72.45 67.17 77.07 78.26 77.43
OlympiadBench Avg@128 14.53 17.94 19.98 21.45 22.31 22.28 22.04 23.67 24.64 24.56
Cons@128 23.41 26.22 27.11 28.00 28.44 29.63 29.48 32.00 33.04 33.78
Pass@64 55.45 55.01 56.70 56.49 55.80 55.84 55.45 56.73 57.57 58.09
MATH-500 Avg@128 39.85 45.48 48.23 50.56 51.41 51.10 49.87 51.22 52.20 52.26
Cons@128 56.80 60.60 62.60 62.00 61.60 62.40 61.80 62.60 65.40 64.60
Pass@64 87.63 87.16 87.38 88.36 88.88 88.81 88.82 89.49 90.18 89.95
Minerva Avg@128 9.11 10.52 11.19 11.99 12.52 12.36 12.93 13.26 12.43 12.75
Cons@128 17.65 17.28 18.38 17.64 19.12 19.49 18.01 20.59 19.12 19.49
Pass@64 44.54 45.07 47.02 44.85 43.97 45.29 48.84 48.53 47.36 47.55
Avg@128 13.90 16.45 17.72 19.06 19.65 19.72 19.11 20.01 20.38 20.49
Average Cons@128 22.56 23.18 24.68 26.41 26.25 27.75 26.83 28.37 28.62 28.26
Pass@64 48.21 48.24 50.44 48.37 48.46 49.35 48.84 50.98 49.83 50.99
Table 19: RL performance of the 4B Dense Model with $\beta=-0.25;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.76 0.96 1.02 0.55 1.09 1.43 1.82 2.03 2.47 2.57
Cons@128 0.00 3.33 3.33 0.00 0.00 3.33 3.33 6.67 6.67 6.67
Pass@64 14.17 17.11 13.31 12.10 17.09 19.99 20.64 19.17 24.12 20.34
AIME25 Avg@128 0.83 1.46 2.11 1.72 2.27 2.11 1.98 2.19 2.40 2.63
Cons@128 0.00 0.00 3.33 3.33 3.33 3.33 3.33 3.33 6.67 3.33
Pass@64 17.46 13.33 21.90 19.12 19.99 15.64 14.15 16.51 11.67 15.82
AMC23 Avg@128 19.55 25.47 29.08 30.96 31.70 31.72 32.38 32.85 33.01 33.09
Cons@128 32.50 40.00 40.00 45.00 45.00 50.00 50.00 45.00 50.00 50.00
Pass@64 72.21 79.29 75.65 74.33 72.62 72.48 70.62 71.24 76.36 74.00
OlympiadBench Avg@128 14.44 17.77 19.71 20.38 21.52 21.44 22.98 23.33 23.52 23.67
Cons@128 24.89 26.52 28.44 28.15 28.59 28.89 31.26 30.81 31.11 31.26
Pass@64 54.46 56.71 57.38 55.02 58.18 55.17 57.97 57.45 56.56 57.23
MATH-500 Avg@128 39.36 45.23 48.29 49.88 50.59 50.72 52.33 52.50 53.09 52.98
Cons@128 58.00 60.40 62.00 60.80 63.20 62.40 62.20 61.60 61.80 62.20
Pass@64 87.37 87.68 88.34 88.53 88.89 89.18 89.02 88.82 89.21 89.94
Minerva Avg@128 9.23 10.71 11.16 12.03 12.14 12.22 12.97 12.73 12.99 13.11
Cons@128 18.39 18.38 18.75 17.65 18.75 18.38 20.96 18.75 18.38 19.49
Pass@64 43.57 46.44 46.34 46.35 46.30 45.61 47.17 45.57 46.61 45.22
Avg@128 14.03 16.93 18.56 19.25 19.89 19.94 20.74 20.94 21.25 21.34
Average Cons@128 22.30 24.77 25.98 25.82 26.48 27.72 28.51 27.69 29.11 28.83
Pass@64 48.21 50.09 50.49 49.24 50.51 49.68 49.93 49.79 50.76 50.43
Table 20: RL performance of the 4B Dense Model with $\beta=0.50;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.94 1.38 1.43 1.46 2.29 1.95 1.74 2.19 2.53 2.42
Cons@128 3.33 3.33 3.33 3.33 6.67 3.33 3.33 3.33 3.33 3.33
Pass@64 18.77 23.78 16.58 20.87 23.65 19.99 18.98 24.56 23.62 22.08
AIME25 Avg@128 0.83 1.88 1.95 2.03 3.39 3.57 2.76 3.67 4.14 4.66
Cons@128 0.00 6.67 3.33 3.33 10.00 10.00 3.33 6.67 10.00 10.00
Pass@64 17.05 15.84 17.51 13.33 15.00 19.59 15.84 20.01 18.77 15.00
AMC23 Avg@128 19.34 22.52 23.09 26.35 26.60 25.46 28.28 29.45 29.49 30.10
Cons@128 32.50 37.50 37.50 40.00 45.00 42.50 37.50 40.00 42.50 42.50
Pass@64 76.20 73.68 75.33 71.56 72.32 74.94 68.79 71.99 73.76 73.35
OlympiadBench Avg@128 14.37 18.07 20.35 21.00 22.38 22.27 22.24 23.32 24.44 24.96
Cons@128 23.56 26.22 28.15 28.59 30.37 29.93 29.33 30.81 31.85 32.00
Pass@64 54.84 57.02 57.95 56.39 57.25 58.01 55.62 57.67 57.93 56.91
MATH-500 Avg@128 39.40 45.70 48.18 49.46 50.41 50.58 51.53 52.32 53.20 54.00
Cons@128 56.00 59.40 61.80 60.20 61.40 61.80 60.20 62.20 62.40 63.20
Pass@64 85.72 87.81 88.00 88.72 88.92 89.49 88.58 88.58 89.39 89.76
Minerva Avg@128 8.50 10.44 11.41 11.34 11.96 11.67 11.73 11.71 12.23 12.10
Cons@128 16.91 16.91 17.65 18.01 18.01 16.91 16.54 15.44 15.44 16.18
Pass@64 42.52 45.32 44.50 43.82 44.10 43.17 41.51 45.40 46.06 43.45
Avg@128 13.90 16.67 17.74 18.61 19.51 19.25 19.71 20.44 21.01 21.37
Average Cons@128 22.05 25.01 25.29 25.58 28.58 27.41 25.04 26.41 27.59 27.87
Pass@64 49.18 50.58 49.98 49.12 50.21 50.87 48.22 51.37 51.59 50.09
Table 21: RL performance of the 4B Dense Model with $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=-0.1$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.70 0.70 0.76 0.89 1.07 0.81 0.86 1.15 1.38 1.46
Cons@128 0.00 3.33 3.33 3.33 3.33 3.33 3.33 3.33 3.33 3.33
Pass@64 22.51 12.51 15.01 14.19 13.77 15.01 20.87 21.03 22.93 23.70
AIME25 Avg@128 0.76 1.09 1.46 1.90 1.67 2.32 2.11 3.02 3.72 3.18
Cons@128 0.00 0.00 0.00 3.33 3.33 6.67 3.33 3.33 6.67 3.33
Pass@64 13.97 15.85 17.50 19.59 18.29 16.68 15.00 21.90 23.58 18.33
AMC23 Avg@128 18.50 20.76 21.84 24.67 21.82 25.35 27.55 28.73 30.47 31.13
Cons@128 37.50 40.00 37.50 40.00 40.00 37.50 37.50 42.50 42.50 45.00
Pass@64 76.21 78.48 79.17 73.11 69.87 77.31 72.07 75.05 73.18 77.07
OlympiadBench Avg@128 13.79 15.29 17.63 19.50 18.03 19.51 20.84 22.23 23.58 24.72
Cons@128 22.22 24.30 25.04 26.81 27.11 27.41 27.41 29.63 29.78 31.70
Pass@64 54.56 56.07 55.14 55.54 56.67 56.07 55.65 56.93 57.90 56.66
MATH-500 Avg@128 37.92 40.85 43.98 45.69 43.78 46.94 48.87 50.44 51.11 53.30
Cons@128 56.20 58.20 59.00 59.00 59.80 59.60 61.80 63.00 64.00 65.00
Pass@64 86.56 87.15 87.23 87.95 87.60 87.54 87.25 88.08 89.22 89.02
Minerva Avg@128 8.07 9.22 10.25 11.08 10.67 11.43 12.28 12.54 12.24 13.13
Cons@128 17.28 18.01 15.07 15.81 17.28 16.18 16.54 18.01 18.38 18.38
Pass@64 41.61 45.66 46.22 43.84 43.71 44.43 44.29 43.08 41.96 42.28
Avg@128 13.29 14.65 15.99 17.29 16.17 17.73 18.75 19.69 20.42 21.25
Average Cons@128 22.20 23.97 23.32 24.71 25.14 25.12 24.99 26.63 27.44 27.79
Pass@64 49.24 49.29 50.05 49.04 48.32 49.51 49.19 51.01 51.46 51.18
Table 22: RL performance of the 4B Dense Model with $\beta=0;\;\tilde{\lambda}=0.1;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.49 0.55 0.91 0.73 0.68 0.63 0.68 0.76 0.91 1.28
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 3.33
Pass@64 9.17 13.24 23.37 8.13 18.97 17.10 20.65 17.94 16.37 20.32
AIME25 Avg@128 0.68 0.83 1.35 1.46 1.28 1.25 1.51 1.93 2.29 2.29
Cons@128 0.00 0.00 3.33 3.33 3.33 3.33 3.33 3.33 3.33 6.67
Pass@64 21.79 22.09 17.83 20.01 16.61 15.81 17.48 15.84 16.26 15.00
AMC23 Avg@128 18.24 20.59 24.16 25.16 23.89 28.38 29.20 29.57 31.43 32.68
Cons@128 32.50 32.50 42.50 35.00 42.50 42.50 42.50 37.50 45.00 45.00
Pass@64 75.30 71.98 75.55 77.70 70.73 71.66 72.38 72.02 75.16 74.30
OlympiadBench Avg@128 13.38 16.45 18.73 19.53 19.28 21.67 21.92 22.44 23.38 24.35
Cons@128 22.96 25.48 27.56 27.70 27.11 29.48 29.33 29.93 31.41 32.00
Pass@64 54.04 55.12 55.80 56.96 56.31 55.42 57.13 56.75 57.94 57.64
MATH-500 Avg@128 37.71 43.26 47.34 48.65 48.09 51.15 51.84 52.49 53.79 54.85
Cons@128 56.00 58.80 62.20 63.20 63.60 63.80 62.80 64.80 64.60 65.40
Pass@64 84.83 87.11 88.42 88.69 87.97 87.80 87.65 88.55 88.98 88.76
Minerva Avg@128 8.00 9.30 10.37 10.83 11.03 11.39 11.47 11.51 11.70 11.66
Cons@128 14.71 16.54 16.91 18.38 16.91 16.91 16.54 18.01 16.54 15.81
Pass@64 41.49 43.44 45.78 44.18 44.17 43.23 44.97 44.22 44.80 45.56
Avg@128 13.08 15.16 17.14 17.73 17.38 19.08 19.44 19.78 20.58 21.19
Average Cons@128 21.03 22.22 25.42 24.60 25.58 26.00 25.75 25.60 26.81 28.04
Pass@64 47.77 48.83 51.13 49.28 49.13 48.50 50.04 49.22 49.92 50.26
Table 23: RL performance of the 20B-A1B MoE Model with $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.34 0.47 0.34 0.44 0.63 0.57 0.55 0.55 0.65 0.65
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 3.33 0.00 0.00
Pass@64 12.88 12.94 11.42 11.68 12.93 11.47 9.59 9.17 10.00 10.84
AIME25 Avg@128 0.13 0.23 0.16 0.29 0.36 0.47 0.60 0.44 0.63 0.57
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Pass@64 6.68 11.27 8.35 14.80 16.58 19.10 18.67 14.10 16.66 16.59
AMC23 Avg@128 10.64 11.82 11.95 13.07 14.26 16.11 16.21 17.79 18.38 19.06
Cons@128 20.00 25.00 17.50 27.50 25.00 30.00 27.50 35.00 32.50 37.50
Pass@64 75.83 76.02 70.67 71.55 73.75 74.33 73.13 73.22 71.65 73.71
OlympiadBench Avg@128 7.13 8.32 9.22 10.66 11.83 12.45 12.60 13.50 14.41 15.03
Cons@128 14.81 16.00 17.04 20.00 21.04 21.33 21.19 22.52 24.30 24.89
Pass@64 49.58 49.53 52.16 52.06 52.69 52.50 53.48 53.34 52.67 53.31
MATH-500 Avg@128 26.71 28.91 32.72 34.81 36.45 37.67 38.47 39.60 41.38 42.19
Cons@128 45.60 48.40 49.80 52.40 54.40 54.40 55.20 56.20 58.20 58.20
Pass@64 82.96 82.71 83.72 83.39 85.93 85.79 85.64 86.01 86.04 85.75
Minerva Avg@128 5.66 5.70 6.42 6.93 7.22 7.84 7.98 8.28 8.36 8.73
Cons@128 9.93 10.66 11.40 12.50 11.40 13.24 12.87 12.50 12.50 13.97
Pass@64 42.39 40.93 41.04 42.75 43.47 42.89 44.35 44.50 40.90 44.90
Avg@128 8.44 9.24 10.14 11.03 11.79 12.52 12.74 13.36 13.97 14.37
Average Cons@128 15.06 16.68 15.96 18.73 18.64 19.83 19.46 21.59 21.25 22.43
Pass@64 45.05 45.57 44.56 46.04 47.56 47.68 47.48 46.72 46.32 47.52
Table 24: RL performance of the 20B-A1B MoE Model with $\beta=-0.25;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.29 0.49 0.86 0.49 0.49 0.86 0.86 0.78 0.86 1.15
Cons@128 0.00 0.00 3.33 0.00 0.00 3.33 0.00 3.33 0.00 3.33
Pass@64 9.15 7.93 8.13 10.84 11.68 14.17 14.59 19.19 15.01 21.27
AIME25 Avg@128 0.16 0.29 0.36 0.36 0.63 0.63 0.89 0.70 0.81 1.04
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Pass@64 8.35 13.78 15.46 7.92 19.81 17.11 19.20 13.12 14.17 17.21
AMC23 Avg@128 11.52 14.02 16.88 18.48 19.47 21.05 22.44 22.66 21.25 22.40
Cons@128 30.00 35.00 37.50 37.50 37.50 32.50 35.00 35.00 32.50 30.00
Pass@64 75.75 73.72 73.54 76.83 74.15 70.35 71.18 74.50 74.13 78.66
OlympiadBench Avg@128 7.08 9.60 11.51 12.79 14.47 15.71 16.85 16.02 16.78 17.68
Cons@128 13.33 17.78 20.59 20.89 21.93 24.59 24.15 24.00 25.19 26.96
Pass@64 47.16 49.12 52.08 51.30 50.62 52.86 53.15 52.78 54.66 55.99
MATH-500 Avg@128 24.43 30.21 34.68 37.65 40.62 42.61 44.71 45.35 44.40 44.86
Cons@128 45.60 48.60 52.60 54.20 54.20 56.20 57.80 59.40 57.00 58.00
Pass@64 81.19 83.88 85.36 84.99 85.75 86.98 88.10 87.39 87.16 86.89
Minerva Avg@128 5.26 6.36 7.45 8.24 9.02 9.89 10.29 10.48 10.14 9.96
Cons@128 10.29 13.24 13.24 13.24 15.07 15.07 15.81 15.81 16.18 16.91
Pass@64 38.70 41.40 43.51 44.46 43.11 45.72 45.94 45.16 45.82 44.46
Avg@128 8.12 10.16 11.96 13.00 14.12 15.13 16.01 16.00 15.71 16.18
Average Cons@128 16.54 19.10 21.21 20.97 21.45 21.95 22.13 22.92 21.81 22.53
Pass@64 43.38 44.97 46.35 46.06 47.52 47.87 48.69 48.69 48.49 50.75
Table 25: RL performance of the 20B-A1B MoE Model with $\beta=0.50;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.42 0.29 0.65 0.70 0.68 0.78 0.68 1.02 0.94 1.04
Cons@128 0.00 0.00 3.33 3.33 0.00 3.33 0.00 3.33 3.33 3.33
Pass@64 18.31 8.32 16.69 19.19 18.77 14.17 12.51 17.53 20.66 16.58
AIME25 Avg@128 0.26 0.16 0.18 0.34 0.26 0.36 0.29 0.26 0.34 0.68
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Pass@64 14.19 9.17 10.01 14.41 13.77 14.97 12.11 11.06 13.67 19.02
AMC23 Avg@128 11.17 13.46 13.54 14.14 15.90 17.27 16.15 17.60 18.42 19.67
Cons@128 17.50 25.00 22.50 25.00 25.00 30.00 27.50 30.00 32.50 32.50
Pass@64 70.56 72.46 71.53 78.57 76.84 78.90 79.26 74.73 78.60 77.41
OlympiadBench Avg@128 7.14 8.17 9.46 10.45 11.60 12.76 12.15 13.38 14.04 15.17
Cons@128 14.81 15.56 16.59 19.11 20.44 21.63 20.44 21.78 22.81 25.19
Pass@64 47.23 50.58 51.54 51.79 52.53 51.78 52.20 52.60 53.27 52.94
MATH-500 Avg@128 27.15 30.51 33.40 34.80 36.67 38.32 36.86 39.42 40.66 41.36
Cons@128 44.40 47.40 47.80 50.60 52.00 52.80 52.20 55.00 55.60 57.00
Pass@64 81.18 82.20 82.85 83.56 84.79 84.78 85.65 86.00 85.97 86.24
Minerva Avg@128 5.86 5.92 6.83 7.01 7.51 8.13 7.94 8.38 8.80 9.01
Cons@128 11.40 10.66 11.76 11.03 12.13 12.13 12.50 13.97 13.60 16.91
Pass@64 41.33 42.59 43.28 42.82 44.07 44.88 44.56 44.13 46.17 45.37
Avg@128 8.67 9.75 10.68 11.24 12.10 12.94 12.35 13.34 13.87 14.49
Average Cons@128 14.69 16.44 17.00 18.18 18.26 19.98 18.77 20.68 21.31 22.49
Pass@64 45.47 44.22 45.98 48.39 48.46 48.25 47.72 47.68 49.72 49.59
Table 26: RL performance of the 20B-A1B MoE Model with $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=-0.1$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.37 0.34 0.50 0.70 0.78 0.70 0.76 0.81 0.78 1.22
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Pass@64 11.45 11.66 13.77 18.58 26.30 17.94 19.83 23.30 26.39 25.86
AIME25 Avg@128 0.39 0.18 0.31 0.31 0.60 0.52 0.55 0.60 0.55 0.76
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Pass@64 12.50 10.84 12.83 13.25 19.91 15.01 23.10 15.03 14.19 13.24
AMC23 Avg@128 11.91 14.30 15.12 17.87 19.71 20.27 21.72 22.68 22.87 24.32
Cons@128 25.00 35.00 27.50 35.00 32.50 37.50 35.00 37.50 35.00 40.00
Pass@64 72.28 75.78 74.15 75.68 77.69 74.47 75.45 70.93 80.24 73.19
OlympiadBench Avg@128 8.08 9.71 10.69 12.87 14.36 14.86 15.96 16.31 16.76 17.47
Cons@128 16.00 20.00 20.89 23.70 23.26 23.26 24.74 25.33 24.59 25.93
Pass@64 49.35 50.50 52.65 54.15 53.54 53.70 54.05 55.85 54.78 54.95
MATH-500 Avg@128 27.41 32.00 34.17 38.08 40.73 41.59 43.06 42.97 44.00 45.20
Cons@128 47.20 51.20 53.60 55.00 57.40 58.00 57.80 58.40 59.40 59.00
Pass@64 84.11 84.77 84.79 86.88 86.96 87.37 86.11 87.32 87.34 87.70
Minerva Avg@128 5.54 6.92 7.44 8.61 9.29 9.17 9.88 9.82 9.92 10.03
Cons@128 11.76 15.44 15.81 16.91 15.44 15.81 16.54 16.91 15.81 18.01
Pass@64 40.01 40.64 41.59 43.33 42.83 44.31 44.58 44.51 44.57 44.25
Avg@128 8.95 10.58 11.37 13.07 14.25 14.52 15.32 15.53 15.81 16.50
Average Cons@128 16.66 20.27 19.63 21.77 21.43 22.43 22.35 23.02 22.47 23.82
Pass@64 44.95 45.70 46.63 48.65 51.21 48.80 50.52 49.49 51.25 49.87
Table 27: RL performance of the 20B-A1B MoE Model with $\beta=0;\;\tilde{\lambda}=0.1;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 0.32 0.50 0.65 0.86 0.83 0.96 0.81 0.68 0.83 0.99
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Pass@64 9.57 10.84 9.17 21.66 10.84 13.96 16.48 10.85 13.67 20.77
AIME25 Avg@128 0.21 0.26 0.34 0.44 0.42 0.76 0.83 0.99 1.17 1.02
Cons@128 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 3.33 3.33
Pass@64 8.97 10.65 14.62 16.81 11.14 18.23 18.34 16.27 17.94 17.51
AMC23 Avg@128 12.30 14.57 17.58 19.16 19.90 20.88 20.55 21.35 20.63 19.86
Cons@128 27.50 30.00 30.00 32.50 30.00 37.5 35.00 40.00 35.00 32.50
Pass@64 71.35 67.79 73.48 73.13 74.08 75.33 77.74 71.64 78.56 76.32
OlympiadBench Avg@128 8.81 11.18 13.36 14.83 15.89 16.77 17.09 16.27 16.30 16.13
Cons@128 15.70 19.11 21.78 23.85 24.59 25.04 24.89 24.44 24.00 24.30
Pass@64 49.50 49.62 50.93 50.98 52.84 52.59 53.86 51.74 51.80 52.61
MATH-500 Avg@128 28.13 33.38 37.19 39.12 41.51 42.88 43.12 43.56 43.53 42.68
Cons@128 48.00 50.40 52.00 53.20 57.00 57.40 57.20 58.40 58.40 58.80
Pass@64 83.25 83.85 86.04 85.39 85.14 85.43 85.67 85.36 86.33 86.73
Minerva Avg@128 5.50 6.58 7.76 8.31 9.08 8.88 8.95 7.83 7.89 7.78
Cons@128 8.46 12.13 11.40 12.50 15.44 15.07 13.97 11.03 9.93 11.40
Pass@64 40.19 40.53 42.63 43.57 44.28 43.39 44.42 42.66 42.90 42.55
Avg@128 9.21 11.08 12.81 13.79 14.61 15.19 15.23 15.11 15.06 14.74
Average Cons@128 16.61 18.61 19.20 20.34 21.17 22.50 21.84 22.31 21.78 21.72
Pass@64 43.81 43.88 46.15 48.59 46.39 48.16 49.42 46.42 48.53 49.42
Table 28: RL performance of the 20B-A1B MoE Model with $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 6.35 8.15 8.98 9.40 11.07 10.55 12.73 13.22 14.01 15.23
Cons@128 20.00 20.00 23.33 23.33 23.30 20.00 23.33 23.33 23.33 23.33
Pass@64 38.19 37.48 34.86 30.84 35.43 37.50 35.00 32.51 30.00 34.59
AIME25 Avg@128 7.81 8.75 8.83 9.69 9.77 10.76 11.22 12.08 12.06 12.97
Cons@128 16.67 16.67 20.00 20.00 16.67 20.00 20.00 16.67 23.33 20.00
Pass@64 89.90 94.10 92.47 92.94 89.63 94.31 92.81 91.25 95.40 94.69
AMC23 Avg@128 36.97 41.91 42.70 43.14 44.34 44.22 46.07 46.60 46.52 48.67
Cons@128 52.50 60.00 55.00 57.50 50.00 52.50 55.00 55.00 55.00 57.50
Pass@64 75.83 76.02 70.67 71.55 73.75 74.33 73.13 73.22 71.65 73.71
OlympiadBench Avg@128 27.07 30.27 31.79 33.02 34.01 35.09 36.67 37.21 37.49 38.67
Cons@128 40.30 41.48 41.93 41.78 42.96 43.26 43.70 44.44 45.04 46.37
Pass@64 69.33 71.45 70.93 70.42 70.64 71.79 71.34 71.61 72.09 71.56
MATH-500 Avg@128 60.32 64.72 66.45 67.54 69.04 69.19 70.95 71.44 71.65 72.67
Cons@128 81.00 81.80 82.60 82.20 82.20 84.20 83.40 84.40 84.80 85.00
Pass@64 96.06 95.76 95.41 95.45 95.69 96.20 96.13 95.83 95.65 95.30
Minerva Avg@128 15.28 17.14 17.67 18.28 19.28 19.05 19.84 19.76 19.80 20.51
Cons@128 26.47 28.68 27.57 27.21 29.41 27.94 28.31 27.94 28.68 30.15
Pass@64 56.03 57.35 56.48 56.99 57.97 57.50 57.52 56.69 57.19 57.63
Avg@128 25.63 28.49 29.40 30.18 31.25 31.48 32.91 33.39 33.59 34.79
Average Cons@128 39.49 41.44 41.74 42.00 40.76 41.32 42.29 41.96 43.36 43.73
Pass@64 64.95 66.36 64.85 64.15 64.97 67.12 65.19 64.51 65.48 65.55
Table 29: RL performance of the 20B-A1B MoE Model with $\beta=-0.25;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 5.08 6.25 7.11 9.51 10.68 10.91 11.67 12.06 12.03 13.38
Cons@128 16.67 20.00 16.67 20.00 23.33 20.00 20.00 20.00 20.00 20.00
Pass@64 35.42 35.22 40.77 42.21 39.60 39.68 40.81 45.54 37.95 46.13
AIME25 Avg@128 34.88 39.82 42.85 45.37 46.80 48.18 49.28 50.45 50.76 52.32
Cons@128 47.50 52.50 55.00 55.00 57.50 60.00 52.50 55.00 55.00 62.50
Pass@64 91.83 93.13 89.55 89.03 89.12 87.25 89.98 92.31 89.64 93.71
AMC23 Avg@128 11.52 14.02 16.88 18.48 19.47 21.05 22.44 22.66 21.25 22.40
Cons@128 30.00 35.00 37.50 37.50 37.50 32.50 35.00 35.00 32.50 30.00
Pass@64 75.75 73.72 73.54 76.83 74.15 70.35 71.18 74.50 74.13 78.66
OlympiadBench Avg@128 24.51 28.36 31.24 33.58 34.83 35.79 37.28 37.75 38.31 39.46
Cons@128 38.07 38.96 42.07 43.41 43.41 44.15 44.89 45.93 46.07 47.70
Pass@64 70.35 70.44 71.28 71.76 71.49 71.44 71.55 71.37 70.84 71.88
MATH-500 Avg@128 57.74 63.71 67.29 69.74 70.84 72.06 73.39 73.85 74.67 75.42
Cons@128 77.40 81.00 82.40 81.80 81.40 82.80 82.80 82.60 82.40 82.80
Pass@64 95.48 95.87 95.50 95.99 95.85 96.12 95.67 95.28 95.39 95.81
Minerva Avg@128 13.98 16.22 18.10 19.01 19.70 20.27 21.41 21.44 21.46 21.88
Cons@128 25.37 26.84 27.57 27.21 26.10 27.94 27.57 27.21 25.74 28.68
Pass@64 55.50 57.87 59.07 58.59 58.57 56.23 57.80 59.11 58.67 57.93
Avg@128 23.66 27.12 29.35 31.22 32.33 33.14 34.26 34.70 35.00 36.06
Average Cons@128 36.39 39.33 39.51 40.68 41.96 41.37 40.74 41.24 41.54 43.06
Pass@64 64.46 64.62 66.40 65.84 65.93 65.33 66.18 67.06 65.35 67.54
Table 30: RL performance of the 20B-A1B MoE Model with $\beta=0.50;\;\tilde{\lambda}=0;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 4.77 5.55 6.64 7.50 7.76 8.20 9.38 9.03 10.18 12.14
Cons@128 13.33 20.00 20.00 20.00 16.67 20.00 20.00 23.33 20.00 23.33
Pass@64 34.76 39.16 36.18 40.52 37.66 38.13 41.45 39.57 39.32 40.99
AIME25 Avg@128 5.83 7.42 7.94 9.61 9.64 9.06 9.74 9.77 10.18 11.25
Cons@128 16.67 20.00 16.67 16.67 16.67 16.67 20.00 23.33 23.33 16.67
Pass@64 32.93 38.56 36.81 39.17 37.86 36.26 37.45 32.88 35.82 42.89
AMC23 Avg@128 30.41 35.27 39.39 40.25 40.66 40.51 43.50 41.13 42.44 45.43
Cons@128 47.50 52.50 60.00 52.50 52.50 52.50 55.00 50.00 55.00 50.00
Pass@64 90.88 93.10 93.04 93.46 93.04 88.67 93.59 91.78 92.50 94.37
OlympiadBench Avg@128 23.16 26.64 27.88 29.90 30.59 31.33 32.83 32.68 33.94 35.50
Cons@128 40.00 42.96 42.96 44.59 44.00 44.15 44.74 45.48 47.56 48.44
Pass@64 70.39 71.35 72.42 72.14 70.88 72.32 71.38 72.65 73.31 73.74
MATH-500 Avg@128 53.63 59.34 60.82 63.07 63.60 64.59 65.94 66.34 66.89 68.51
Cons@128 79.20 81.60 81.20 82.20 81.60 81.40 83.40 84.00 83.20 84.20
Pass@64 95.45 96.13 96.11 96.92 95.88 95.08 96.79 95.69 96.29 96.88
Minerva Avg@128 14.43 16.70 18.01 19.24 19.58 20.07 20.61 20.83 20.91 21.60
Cons@128 25.74 29.41 28.31 29.41 29.78 30.88 30.88 30.51 30.88 30.51
Pass@64 55.45 59.42 59.01 58.72 58.51 58.87 59.29 59.13 59.93 59.30
Avg@128 22.04 25.15 26.78 28.26 28.64 28.96 30.33 29.96 30.76 32.41
Average Cons@128 37.07 41.08 41.52 40.90 40.20 40.93 42.34 42.78 43.33 42.19
Pass@64 63.31 66.29 65.60 66.82 65.64 64.89 66.66 65.28 66.20 68.03
Table 31: RL performance of the 20B-A1B MoE Model with $\beta=0;\;\tilde{\lambda}=0;\;\hat{\lambda}=-0.1$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 6.72 8.67 9.77 11.85 11.12 12.29 12.81 13.49 14.14 15.03
Cons@128 20.00 20.00 20.00 23.33 23.33 20.00 20.00 23.33 20.00 23.33
Pass@64 38.01 42.74 35.32 38.36 38.76 39.20 40.77 41.44 40.38 43.95
AIME25 Avg@128 7.97 9.58 10.39 10.52 11.64 11.77 12.47 12.66 13.62 13.98
Cons@128 13.33 16.67 13.33 23.33 23.33 23.33 23.33 23.33 23.33 23.33
Pass@64 37.42 42.73 39.13 41.91 41.47 35.01 37.31 45.65 41.49 42.49
AMC23 Avg@128 36.45 40.68 42.64 45.55 45.29 46.68 49.82 49.71 50.70 51.29
Cons@128 52.50 57.50 52.50 52.50 57.50 57.50 60.00 60.00 60.00 62.50
Pass@64 94.89 95.79 93.11 93.75 94.31 93.25 92.25 91.06 91.75 92.45
OlympiadBench Avg@128 27.17 30.22 32.75 34.63 35.71 35.94 37.33 38.01 38.85 39.91
Cons@128 40.15 43.11 45.04 46.22 46.81 45.04 46.67 47.85 48.00 49.48
Pass@64 70.80 72.13 72.46 71,75 71.95 71.81 72.33 71.94 72.48 72.09
MATH-500 Avg@128 60.69 64.29 67.51 69.76 71.15 72.00 73.20 73.95 75.17 74.46
Cons@128 81.00 82.20 82.20 82.80 83.40 83.80 84.00 84.80 84.80 84.80
Pass@64 95.94 96.53 96.13 95.96 96.18 95.47 95.62 96.21 96.24 96.26
Minerva Avg@128 13.84 15.75 16.99 18.44 19.28 20.27 20.62 20.63 21.20 21.47
Cons@128 24.26 25.37 25.00 25.37 27.94 27.21 28.68 28.68 29.41 29.41
Pass@64 57.22 58.83 58.68 59.09 60.42 58.27 58.51 58.19 58.62 58.18
Avg@128 25.47 28.20 30.01 31.79 32.37 33.16 34.38 34.74 35.61 36.02
Average Cons@128 38.54 40.81 39.68 42.26 43.72 42.81 43.78 44.67 44.26 45.48
Pass@64 65.71 68.13 65.81 65.81 67.18 65.50 66.13 67.42 66.83 67.57
Table 32: RL performance of the 20B-A1B MoE Model with $\beta=0;\;\tilde{\lambda}=0.1;\;\hat{\lambda}=0$.
# RL Steps 100 200 300 400 500 600 700 800 900 1000
AIME24 Avg@128 5.76 6.77 8.33 9.27 9.11 9.89 10.21 10.13 11.07 11.93
Cons@128 13.33 13.33 23.33 20.00 20.00 23.33 23.33 20.00 20.00 23.33
Pass@64 37.39 32.44 33.13 35.41 31.26 37.10 30.85 33.15 31.27 30.43
AIME25 Avg@128 8.02 9.09 9.77 11.28 11.20 11.35 11.07 11.82 12.06 12.19
Cons@128 13.33 16.67 16.67 23.33 16.67 16.67 16.67 20.00 16.67 20.00
Pass@64 37.39 38.82 39.48 39.60 40.46 34.03 39.78 34.90 39.92 38.34
AMC23 Avg@128 37.64 41.50 42.64 45.20 46.04 47.40 45.72 46.13 47.93 48.59
Cons@128 55.00 55.00 52.50 55.00 52.50 57.50 52.50 52.50 60.00 52.50
Pass@64 90.88 93.57 90.15 93.91 90.47 92.18 94.23 89.84 92.39 94.66
OlympiadBench Avg@128 27.11 30.38 32.26 33.98 34.79 35.56 34.90 35.84 36.77 37.59
Cons@128 40.30 41.48 42.07 43.85 43.70 44.00 43.56 44.30 44.74 45.93
Pass@64 70.82 71.19 70.78 71.88 71.05 70.06 70.52 70.86 71.53 71.13
MATH-500 Avg@128 60.96 65.60 68.09 60.68 70.52 71.22 71.23 71.75 72.69 73.29
Cons@128 79.60 81.00 81.80 82.40 80.80 82.00 81.80 82.00 82.60 82.40
Pass@64 95.66 96.17 95.63 96.09 95.93 95.57 95.12 94.99 95.67 95.53
Minerva Avg@128 13.85 16.13 17.85 18.91 18.98 19.84 19.78 19.82 20.69 21.17
Cons@128 25.00 26.84 27.94 28.31 27.94 28.31 28.31 25.37 29.04 27.57
Pass@64 55.91 55.65 56.66 56.81 57.18 58.14 56.76 57.47 57.85 57.34
Avg@128 25.56 28.24 29.82 29.89 31.77 32.54 32.15 32.58 33.54 34.13
Average Cons@128 37.76 39.05 40.72 42.15 40.27 41.97 41.03 40.70 42.18 41.96
Pass@64 64.68 64.64 64.31 65.62 64.39 64.51 64.54 63.54 64.77 64.57